### v1

# A Unified Voxel Diffusion Module for Point Cloud 3D Object Detection

Abstract
Recent advances in point cloud object detection have increasingly adopted Transformer-based and State Space Models (SSMs), demonstrating strong performance. However, voxel-based representations in these models require strict consistency in input and output dimensions due to their serialized processing, which limits the spatial diffusion capability typically offered by convolutional operations. This limitation significantly affects detection accuracy. Inspired by CNN-based object detection architectures, we propose a novel Voxel Diffusion Module (VDM) to enhance voxel-level representation and diffusion in point cloud data. VDM is composed of sparse 3D convolutions, submanifold sparse convolutions, and residual connections. To ensure computational efficiency, the output feature maps are downsampled to one-fourth of the original input resolution. VDM serves two primary functions: (1) diffusing foreground voxel features through sparse 3D convolutions to enrich spatial context, and (2) aggregating fine-grained spatial information to strengthen voxel-wise feature representation. The enhanced voxel features produced by VDM can be seamlessly integrated into mainstream Transformer- or SSM-based detection models for accurate object classification and localization, highlighting the generalizability of our method. We evaluate VDM on several benchmark datasets by embedding it into both Transformer-based and SSM-based models. Experimental results show that our approach consistently improves detection accuracy over baseline models. Specifically, VDM-SSMs achieve 74.7 mAPH (L2) on Waymo, 72.9 NDS on nuScenes, 42.3 mAP on Argoverse 2, and 67.6 mAP on ONCE, setting new state-of-the-art performance across all datasets. Our code will be made publicly available.
Introduction
3D object detection based on point clouds plays a crucial role in applications such as autonomous driving (Bansal, Krizhevsky, and Ogale 2018) and robotic navigation (Zhu et al. 2017). Inspired by the remarkable success of Transformers (Vaswani et al. 2017) and State Space Models (SSMs) (Gu and Dao 2023) in natural language processing, researchers have recently explored adapting these architectures to computer vision tasks, achieving strong performance. For example, in the domain of point cloud 3D object detection, DVST (Wang et al. 2023) introduces a dynamic sparse window attention mechanism that adaptively partitions each window into local sub-regions based on voxel sparsity, enabling fully parallel attention computations. LION (Liu et al. 2024) further extends this direction by employing large grouping windows and leveraging SSMs’ ability to capture long-range voxel dependencies, setting new state-of-the-art (SOTA) performance on several public datasets.
Figure 1: Performance comparison of different models on various datasets. Numbers highlighted in red indicate that our model achieves state-of-the-art results.
Despite these advancements, both Transformer-based and SSM-based models for point cloud voxel sequence processing require consistency in input and output sequence dimensions, which fundamentally limits their ability to perform convolution-like spatial diffusion. However, prior work (Zhang et al. 2024a; Liu et al. 2024) has demonstrated that voxel diffusion, particularly for foreground voxels, is critical for enhancing detection accuracy. For instance, LION selects top-$k$ voxels via feature summation and applies directional diffusion during intermediate processing. In contrast, our method proposes to conduct voxel diffusion before serialization, using convolutional operations to enrich the original voxel space with contextually relevant features.
Motivated by CNN-based object detection architectures, we propose a novel Voxel Diffusion Module (VDM) tailored for both Mamba-based and Transformer-based 3D detection models. Positioned before voxel sequence serialization, VDM leverages 3D convolutional operations to diffuse foreground voxel features and increase their spatial density. Simultaneously, sparse convolutions are employed to aggregate fine-grained spatial features within each voxel’s local 3D neighborhood, supplying the downstream sequence model with richer spatial context. Structurally, VDM consists of stacked sparse 3D convolutions and residual blocks, effectively supporting both diffusion and aggregation. The output feature map is downsampled to one-fourth of the input resolution (in height and width), helping reduce the computational load for subsequent models.
VDM is designed to be modular and general, allowing seamless integration with both Transformer-based and SSM-based detectors. To evaluate its effectiveness, we perform extensive experiments on four large-scale datasets. As shown in Figure 1, our VDM-enhanced SSM-based model (VDM-Mamba) achieves 74.7 mAPH (L2) on Waymo, 68.1 mAP on nuScenes, 42.3 mAP on Argoverse 2, and 67.6 mAP on ONCE, setting a new state-of-the-art (SOTA) performance. Additionally, the Transformer-based variant (VDM-DSVT) improves performance by +2.2 mAPH (L2) over the DSVT baseline on the Waymo dataset, confirming the general effectiveness of our approach.
Our contributions can be summarized as follows:
  • We propose VDM, a voxel diffusion module that enhances the representation of raw voxel inputs through 3D convolution, significantly enriching foreground voxel features and improving detection accuracy.
  • VDM utilizes sparse 3D convolutions and residual networks to aggregate spatial features from the voxel neighborhood, providing richer positional cues for downstream sequence modeling.
  • VDM is a general-purpose module that integrates well with both Transformer-based and SSM-based 3D detectors, consistently boosting model performance.
  • We validate the effectiveness of VDM through experiments on multiple public datasets, where VDM-SSMs achieve SOTA results, and VDM also improves Transformer-based models.
Related Work
3D object detection based on point clouds has seen significant progress in recent years. According to model architecture, existing approaches can be broadly categorized into three types: CNN-based, Transformer-based, and SSM-based methods. CNN-based models emerged earlier and have been thoroughly studied, achieving solid performance. In contrast, Transformer and Mamba-based models are relatively recent, and while less explored for 3D object detection, they have already shown promising results—often surpassing CNN-based methods in detection accuracy.
CNN-Based 3D Object Detection. CNN-based 3D detectors are typically categorized into voxel-based and point-based approaches. VoxelNet (Zhou and Tuzel 2018) introduces a Voxel Feature Encoding (VFE) layer to convert point clusters within voxels into unified feature representations. HEDNet uses an encoder-decoder design to capture long-range spatial dependencies, particularly useful for detecting large or distant objects. SAFDNet proposes an adaptive feature diffusion strategy to address missing center features in sparse detectors. PointPillars (Lang et al. 2019) leverages PointNet to learn features from vertical columns of point clouds, offering notable inference speed advantages. CenterPoint (Yin, Zhou, and Krahenbuhl 2021) detects object centers using a keypoint detector and regresses attributes such as 3D size, orientation, and velocity. 3DSSD (Yang et al. 2020) introduces a fused sampling strategy in the downsampling process to retain informative representative points. IASSD (Zhang et al. 2022) proposes two task-aware, instance-sensitive sampling strategies for hierarchically selecting foreground points of interest. Among these approaches, voxel-based detectors generally achieve higher accuracy and have become the dominant choice for outdoor 3D detection tasks.
Transformer-Based 3D Object Detection. Transformers have achieved state-of-the-art performance in both NLP and computer vision tasks. PVT-SSD (Yang et al. 2023) introduces a Point-Voxel Transformer that captures long-range dependencies via voxel tokens while preserving fine spatial detail through point-level features. FlatFormer (Liu et al. 2023) flattens point clouds using window-based sorting and partitions them into equal-sized groups, improving detection efficiency. DSVT (Wang et al. 2023) proposes a Dynamic Sparse Window Attention mechanism that adaptively splits each window based on voxel sparsity, enabling fully parallel feature computation. However, the serialization of voxel features in Transformer-based methods can lead to the loss of fine-grained spatial cues, limiting performance in complex 3D scenes.
SSM-Based 3D Object Detection. SSM-based detectors represent a new and emerging line of work. LION (Liu et al. 2024) applies a large-group strategy during voxel serialization for more effective spatial context modeling. Voxel Mamba (Zhang et al. 2024b) eliminates the grouping step and serializes the entire voxel space as a single sequence, introducing a Dual-Scale SSM Block to capture hierarchical spatial dependencies. UniMamba combines 3D convolution with SSMs in a compact multi-head architecture that efficiently captures both local and global spatial context.
Although still in their early stages, SSM-based models already outperform traditional CNN-based frameworks on several benchmarks due to their strong modeling capabilities. Our proposed Voxel Diffusion Module (VDM) further enhances these serialized architectures by introducing voxel diffusion and fine-grained spatial feature aggregation prior to serialization. This design supplies Transformer- and SSM-based models with richer input representations, leading to significant performance gains in 3D object detection.
Figure 2: Overall architecture of our proposed method. (a) The Voxel Diffusion Module (VDM) is placed before the sequential detection model, enabling diffusion of raw voxelized point clouds and aggregation of spatial voxel features. Internally, it consists of Submanifold 3D Convolution (SubM3D), Sparse Residual Block (SRB-3D), and Sparse 3D Convolution (SPConv-3D), where SPConv-3D plays a key role in voxel diffusion. (b) Mamba Block. (c) Sparse Residual Block (SRB-3D). The red connection line ”Only-Diffusion” in the figure indicates the model architecture when only the voxel diffusion strategy is applied in the VDM, aiming to highlight the critical impact of voxel diffusion on detection performance.
Method
In this section, we introduce the proposed Voxel Diffusion Module (VDM), designed to enhance the representation of voxelized point cloud features. VDM leverages convolutional operations to perform early-stage voxel diffusion, increasing the density of meaningful voxel information. This provides richer input for downstream serialized models such as Transformers or Mamba. Furthermore, VDM aggregates local spatial features within each voxel’s 3D neighborhood, enabling the sequence models to more effectively utilize spatial positional cues during processing.
Preliminaries
Transformer. Transformers, based on self-attention mechanisms, have been widely adopted in both natural language processing and computer vision. The standard attention mechanism is defined as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (1)
Here, $Q$ (Query), $K$ (Key), and $V$ (Value) represent the input projection vectors. $Q$ interacts with $K^{\top}$ to compute similarity scores, which are then used to weight the values in $V$. The term $d_{k}$ denotes the dimensionality of the key vectors, and the scaling factor $\sqrt{d_{k}}$ is introduced to stabilize the gradients during training. While Transformers are powerful in capturing global dependencies, their quadratic time complexity poses a challenge when handling long sequences, such as in voxelized point clouds.
State Space Models (SSMs). SSMs model sequential data through continuous-time state equations:
$h^{\prime}(t)=Ah(t)+Bx(t)$ (2)
$y(t)=Ch(t)$ (3)
where $x(t)\in\mathbb{R}^{L}$ is the continuous input signal, $y(t)\in\mathbb{R}^{L}$ is the output, and $A\in\mathbb{R}^{N\times N}$, $B\in\mathbb{R}^{N\times 1}$, and $C\in\mathbb{R}^{1\times N}$ are learnable parameters. The discrete form is expressed as:
$h_{t}=A^{\prime}h_{t-1}+B^{\prime}x_{t}$ (4)
$y_{t}=Ch_{t}$ (5)
where $A^{\prime}$ and $B^{\prime}$ are obtained from continuous $A$ and $B$ via a discretization rule: $A^{\prime}=f_{A}(\Delta,A)$ and $B^{\prime}=f_{B}(\Delta,A,B)$. Mamba-based SSM models (Gu and Dao 2023) have demonstrated significant advantages over Transformers in terms of modeling long-range dependencies and computational efficiency. In the context of 3D point cloud detection, such models (Jin et al. 2025; Liu et al. 2024) have shown promising results and even outperformed many Transformer-based architectures.
Sparse Residual Block (SRB). In CNN-based 3D detectors such as SAFDNet (Zhang et al. 2024a), the Sparse Residual Block (SRB) is commonly used for efficient feature extraction. An SRB consists of two 3D submanifold sparse convolution layers and a residual skip connection linking the input to the output. In our design, VDM is composed of stacked SRB modules and additional SubM3D layers to enable both voxel diffusion and spatial feature aggregation. This structure facilitates the enrichment of voxel-level spatial representations before they are passed into the serialized backbone (Transformer or SSM), thereby improving detection accuracy.
Figure 3: Visualization of the Voxel Diffusion Module (VDM) applied to the original point clouds, demonstrating its effect in voxel-level diffusion and feature aggregation.
Overall Architecture
As illustrated in Figure 2, the Voxel Diffusion Module (VDM) is designed to perform voxel-level diffusion of point clouds and aggregate fine-grained voxel features. This facilitates the generation of rich foreground voxel representations and fine-grained spatial features, serving as a beneficial foundation for subsequent transformer-based and state-space model (SSM)-based detectors. Convolutional neural network (CNN)-based models (Zhang et al. 2024a) have been extensively studied in 3D object detection. These models typically begin by applying feature extraction networks to voxelized point clouds, followed by a backbone and a detection head for final predictions. Inspired by this paradigm, the proposed VDM is composed of alternating Submanifold 3D Convolution (SubM3D) layers and Sparse Residual Blocks (SRB). Additionally, two Sparse 3D Convolution (SPConv3D) layers are inserted to downsample and upsample the voxel feature maps, enabling hierarchical feature learning and effective diffusion.
Voxel Diffusion in VDM. Voxel diffusion plays a significant role in enhancing model detection performance, as demonstrated in previous studies (Liu et al. 2024; Zhang et al. 2024a). Inspired by the property of convolution operations to increase the number of sparse features, we introduce a sparse 3D convolution with a kernel size of 3 within our proposed Voxel Diffusion Module (VDM). This module enables diffusion over the input voxel grid, thus expanding the spatial distribution of informative voxels. Figure 3 illustrates how VDM diffuses the original point cloud into a denser voxel representation, providing a richer feature basis for subsequent serial detection networks. To quantitatively evaluate this effect, we perform a statistical analysis on the voxel count post-VDM processing, as shown in Table 1. As observed, the number of total and foreground voxels significantly increases after applying VDM. This expansion ensures that subsequent modules—particularly those based on transformers or State Space Models (SSMs)—can access more comprehensive foreground point features, thereby improving object detection accuracy. It is worth noting that voxel diffusion serves as a crucial component of the VDM. Our later experiments will further highlight its contribution to performance gains, underlining its importance in boosting the effectiveness of serialization-based architectures such as transformers and Mamba variants.
Figure 4: Fine-grained feature aggregation within the VDM, utilizing convolutional operations to aggregate relative spatial information across voxels.
Method Overall Voxel Count Foreground Voxel Count mAPH L2
Non(LION) 36340 2616 74.0
VDM 188562 9146 74.8
VDM(OD) 216601 10705 74.8
Table 1: Effect of VDM on voxel counts in input point clouds on the Waymo validation dataset. The overall and foreground voxel counts are computed as the total number of corresponding voxels across all frames divided by the total number of frames.
Methods Operator Vehicle 3D AP/APH Pedestrian 3D AP/APH Cyclist 3D AP/APH mAP/mAPH (L2)
L1 L2 L1 L2 L1 L2
SECOND
CNN
72.3/71.7 63.9/63.3 68.7/58.2 60.7/51.3 60.6/59.3 58.3/57.0 61.0/57.2
PointPillars 72.1/71.5 63.6/63.1 70.6/56.7 62.8/50.3 64.4/62.3 61.9/59.9 62.8/57.8
CenterPoint 74.2/73.6 66.2/65.7 76.6/70.5 68.8/63.2 72.3/71.1 69.7/68.5 68.2/65.8
PV-RCNN 78.0/77.5 69.4/69.0 79.2/73.0 70.4/64.7 71.5/70.3 69.0/67.8 69.6/67.2
PillarNet-18 78.2/77.7 70.4/69.9 79.8/72.6 71.6/64.9 70.4/69.3 67.8/66.7 69.9/67.2
FSD 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3 72.9/70.8
AFDetV2 77.6/77.1 69.7/69.2 80.2/74.6 72.2/67.0 73.7/72.7 71.0/70.1 71.0/68.8
PillarNeXt 78.4/77.9 70.3/69.8 82.5/77.1 74.9/69.8 73.2/72.2 70.6/69.6 71.9/69.7
VoxelNext 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2 72.2/70.1
CenterFormer 75.0/74.4 69.9/69.4 78.6/73.0 73.6/68.3 72.3/71.3 69.8/68.8 71.1/68.9
PV-RCNN++ 79.3/78.8 70.6/70.2 81.3/76.3 73.2/68.0 73.7/72.7 71.2/70.2 71.7/69.5
ConQueR 76.1/75.6 68.7/68.2 79.0/72.3 70.9/64.7 73.9/72.5 71.4/70.1 70.3/67.7
FocalFormer3D –/– 68.1/67.6 –/– 72.7/66.8 –/– 73.7/72.6 71.5/69.0
HEDNet 81.1/80.6 73.2/72.7 84.4/80.0 76.8/72.6 78.7/77.7 75.8/74.9 75.3/73.4
VPF 80.2/79.7 71.9/71.5 82.5/76.9 74.8/69.4 77.1/76.0 74.2/73.2 73.6/71.4
TransFusion
Transformer
–/– –/65.1 –/– –/63.7 –/– –/65.9 –/64.9
SST_TS 76.2/75.8 68.0/67.6 81.4/74.0 72.8/65.9 –/– –/– –/–
SWFormer 77.8/77.3 69.2/68.8 80.9/72.7 72.5/64.9 –/– –/– –/–
FlatFormer - 69.0/68.6 - 71.5/65.3 –/– 68.6/67.5 69.7/67.2
OcTr 78.1/77.6 69.8/69.3 80.8/74.4 72.5/66.5 72.6/71.5 69.9/68.9 70.7/68.2
DSVT-Pillar 79.3/78.8 70.9/70.5 82.8/77.0 75.2/69.8 76.4/75.4 73.6/72.7 73.2/71.0
DSVT-Voxel 79.2/78.8 70.9/70.5 82.9/77.7 75.2/70.3 77.2/76.2 74.3/73.4 73.5/71.4
VDM-DSVT(Ours) 79.7/79.3 71.4/71.0 84.3/79.7 76.7/72.3 78.4/77.4 75.6/74.6 74.6/72.6
Voxel Mamba
SSMs
80.8/80.3 72.6/72.2 85.0/80.8 77.7/73.6 78.6/77.6 75.7/74.8 75.3/73.6
LION 80.3/79.9 72.0/71.6 85.8/81.4 78.5/74.3 80.1/79.0 77.2/76.2 75.9/74.0
UniMamba 80.6/80.1 72.3/71.8 86.0/81.3 78.7/74.1 80.3/79.3 77.5/76.5 76.1/74.1
VDM-Mamba (Ours) 81.0/80.5 72.6/72.2 86.4/82.5 79.3/75.5 80.4/79.4 77.5/76.5 76.5/74.8
VDM-OD-Mamba (Ours) 81.3/80.9 73.0/72.6 86.1/82.1 78.9/75.0 80.7/79.7 77.8/76.9 76.6/74.8
Table 2: Comparison with prior methods (CNNs, Transformers, and SSMs) on the Waymo Open Dataset using 100% training data. Metrics are reported as mAP/mAPH (%) $\uparrow$ for overall results and AP/APH (%) $\uparrow$ per category. All models are trained under the single-frame setting. “–” indicates unavailable results. VDM-OD-Mamba denotes a variant where the Only-Diffusion structure is used within the VDM framework. The best-performing results are highlighted in bold.
Fine-Grained Feature Aggregation. To capture more detailed spatial information, we further divide the voxel grid into finer partitions, as illustrated in Figure 4. Leveraging the strong spatial perception capabilities of the Submanifold 3D Convolution (SubM3D) and the Sparse Residual Block (SRB) within the VDM, we extract fine-grained voxel features from these subdivided regions. In addition, we adopt a sparse convolution with a stride of 2 to serve dual purposes: first, it reduces the resolution of the feature map to half, which effectively controls computational cost by mitigating the overhead introduced by voxel diffusion; second, it aggregates local fine-grained features, thereby enriching the spatial context available for downstream modules. This design ensures that the model benefits from both detailed spatial representations and efficient computation during object detection.
The function of VDM is to diffuse the number of sparse voxels and to aggregate fine-grained features. It can be summarized by the following formulas:
$\mathit{FG}^{\prime}=\mathrm{Voxel\_Diffusion}(\mathit{FG},\mathit{spconv})$ (6)
$\mathit{F}^{\prime}=\mathrm{Aggregation}(\mathit{F},\mathit{subm},\mathit{spconv})$ (7)
where $\mathit{F}$ denotes the initial voxel feature map; $\mathit{F}^{\prime}$ represents the feature map after local voxel aggregation, with width and height reduced to one-fourth of the original size; $\mathit{subm}$ denotes submanifold sparse convolution, and $\mathit{spconv}$ denotes sparse convolution. $\mathit{FG}^{\prime}$ indicates the number of foreground voxels after VDM processing, and $\mathit{FG}$ is the number of foreground voxels in the initial feature map.
Integration with SSMs and Transformer Modules. The proposed VDM module can be integrated with both Mamba-based and Transformer-based models. By effectively diffusing sparse voxels and aggregating fine-grained point cloud voxel features, VDM further enhances the detection accuracy of the overall model. For the 3D feature maps processed by VDM, we first serialize them into one-dimensional sequences along the $x$ and $y$ directions. As the sequence length increases, Mamba-based networks are able to adopt larger grouping sizes due to their lower computational complexity compared to Transformer-based counterparts. Given the strong performance of LION (Liu et al. 2024) and DSVT (Wang et al. 2023) in the field of 3D object detection, we integrate the VDM module into both networks and conduct experiments to validate its effectiveness.
Experiments
Datasets and Evaluation Metrics
Waymo Open Dataset. The Waymo Open Dataset (WOD) (Sun et al. 2020) includes 1,150 scenes, divided into 798 for training, 202 for validation, and 150 for testing. Each scene covers a perception range of $150\,\text{m}\times 150\,\text{m}$. For evaluation, WOD adopts 3D mean Average Precision (mAP) and mAP weighted by heading accuracy (mAPH). Each metric includes two difficulty levels: L1 for objects with more than five points, and L2 for bounding boxes containing one to five points.
nuScenes Dataset. The nuScenes dataset (Caesar et al. 2020) contains 1,000 scenes, with 750 for training, 150 for validation, and 150 for testing. It uses mean Average Precision (mAP) and nuScenes Detection Score (NDS) as evaluation metrics. The dataset covers 10 annotated object classes, representing most common categories in traffic scenarios.
Argoverse 2 Dataset. The Argoverse 2 dataset (Chang et al. 2019) comprises 1,000 sequences, with 700 for training, 150 for validation, and 150 for testing. The perception range extends up to 200 meters. It uses mean Average Precision (mAP) as the evaluation metric.
ONCE Dataset. The ONCE dataset (Mao et al. 2021) consists of 5,000, 3,000, and 8,000 frames for training, validation, and testing respectively. It defines three detection classes: Vehicle, Pedestrian, and Cyclist. The final performance is evaluated using mean Average Precision (mAP) across all three classes.
Performances on the validation set
Method NDS mAP Car Truck Bus T.L. C.V. Ped. M.T. Bike T.C. B.R.
CenterPoint 66.5 59.2 84.9 57.4 70.7 38.1 16.9 85.1 59.0 42.0 69.8 68.3
VoxelNeXt 66.7 60.5 83.9 55.5 70.5 38.1 21.1 84.6 62.8 50.0 69.4 69.4
Uni3DETR 68.5 61.7
Transfusion-LiDAR 70.1 65.5 86.9 60.8 73.1 43.4 25.2 87.5 72.9 57.3 77.2 70.3
DSVT 71.1 66.4 87.4 62.6 75.9 42.1 25.3 88.2 74.8 58.7 77.9 71.0
HEDNet 71.4 66.7 87.7 60.6 77.8 50.7 28.9 87.1 74.3 56.8 76.3 66.9
SAFDNet 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
LION 72.1 68.0 87.9 64.9 77.6 44.4 28.5 89.6 75.6 59.4 80.8 71.6
UniMamba 72.6 68.5 88.7 64.7 79.7 47.9 28.7 89.7 74.6 59.1 79.5 72.3
VDM-mamba(Ours) 72.5 68.1 88.5 59.5 78.8 43.2 30.6 90.2 75.2 61.1 81.4 72.5
VDM-OD-mamba(Ours) 72.9 68.5 88.7 65.7 78.9 43.8 29.7 90.2 75.7 60.9 81.1 70.3
Performances on the test set
TransFusion-LiDAR 70.2 65.5 86.2 56.7 66.3 58.8 28.2 86.1 68.3 44.2 82.0 78.2
DSVT 72.7 68.4 86.8 58.4 67.3 63.1 37.1 88.0 73.0 47.2 84.9 78.4
HEDNet 72.0 67.7 87.1 56.5 70.4 63.5 33.6 87.9 70.4 44.8 85.1 78.1
VPF 72.7 67.0 85.8 55.1 63.5 62.1 33.3 87.6 72.5 48.6 82.9 78.2
SAFDNet 72.3 68.3 87.3 57.3 68.0 63.7 37.3 89.0 71.1 44.8 84.9 79.5
LION 73.9 69.8 87.2 61.1 68.9 65.0 36.3 90.0 74.0 49.2 87.3 79.5
VDM-mamba(Ours) 73.7 70.0 88.0 60.1 67.6 66.0 35.1 90.9 72.2 50.9 89.2 80.0
Table 3: Performances on the nuScenes validation and test set. ‘T.L.’, ‘C.V.’, ‘Ped.’, ‘M.T.’, ‘T.C.’, and ’B.R.’ are short for trailer, construction vehicle, pedestrian, motor, traffic cone, and barrier, respectively. VDM-OD-Mamba denotes a variant where the Only-Diffusion structure is used within the VDM framework.
Method
mAP
Vehicle
Bus
Pedestrian
Stop Sign
Box Truck
Bollard
C-Barrel
Motorcyclist
MPC-Sign
Motorcycle
Bicycle
A-Bus
Shool Bus
Truck Cab
C-Cone
V-Trailer
Sign
Large Vehicle
Stroller
Bicyclist
Truck
MBT
Dog
Wheelchair
W-Device
W-Rider
CenterPoint 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1 22.1 0.0 3.9 0.5 10.9 4.2
HEDNet 37.1 78.2 47.7 67.6 46.4 45.9 56.9 67.0 48.7 46.5 58.2 47.5 23.3 40.9 27.5 46.8 27.9 20.6 6.9 27.2 38.7 21.6 0.0 30.7 9.5 28.5 8.7
VoxelNeXt 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
FSDV2 37.6 77.0 47.6 70.5 43.6 41.5 53.9 58.5 56.8 39.0 60.7 49.4 28.4 41.9 30.2 44.9 33.4 16.6 7.3 32.5 45.9 24.0 1.0 12.6 17.1 26.3 17.2
SAFDNet 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
LION 41.5 75.1 43.6 73.9 53.9 45.1 66.4 74.7 61.3 48.7 65.1 56.2 21.7 42.7 25.3 58.4 28.9 23.6 8.3 49.5 47.3 19.0 0.0 31.4 8.7 37.6 11.8
UniMamba 42.0 78.9 47.9 74.3 51.8 46.8 67.8 76.9 55.8 51.7 62.8 52.2 30.2 44.6 24.6 28.1 59.4 32.2 23.2 6.7 41.5 48.5 0.0 26.4 8.1 36.4 13.7
VDM (Ours) 42.3 78.2 48.2 75.2 53.2 45.7 68.2 77.4 60.3 47.2 67.0 58.0 24.9 43.0 24.9 60.1 31.4 25.1 7.8 42.5 46.3 20.9 0.0 29.7 10.3 39.8 13.6
VDM-OD(Ours) 42.6 79.0 49.5 75.9 54.5 45.5 69.9 77.9 56.9 47.6 66.6 59.2 24.0 44.2 24.7 60.9 33.2 24.9 7.5 39.5 51.3 20.1 0.0 33.7 7.8 40.1 12.5
Table 4: Comparison with prior methods on the Argoverse2 validation set. ‘Vehicle’, ‘C-Barrel’, ‘MPC-Sign’, ‘A-Bus’, ‘C-Cone’, ‘V-Trailer’, ‘MBT’, ‘W-Device’, and ‘W-Rider’ are abbreviations for regular vehicle, construction barrel, mobile pedestrian crossing sign, articulated bus, construction cone, vehicular trailer, message board trailer, wheeled device, and wheeled rider, respectively. VDM and VDM-OD refer to the proposed VDM-Mamba and VDM-OD-Mamba variants.
Implementation Details
Network Architecture. We evaluate the proposed VDM module by integrating it with both the SSMs-based LION (Liu et al. 2024) and the Transformer-based DSVT (Wang et al. 2023) models. We use voxel sizes of $(0.075\,\text{m},0.075\,\text{m},0.25\,\text{m})$, $(0.08\,\text{m},0.08\,\text{m},0.1875\,\text{m})$, $(0.1\,\text{m},0.1\,\text{m},0.25\,\text{m})$, and $(0.1\,\text{m},0.1\,\text{m},0.25\,\text{m})$ for the nuScenes, Waymo, Argoverse 2, and ONCE datasets respectively. The number of voxel feature channels is set to 64.
In the VDM module, the SubM3D (including SRB) layers use a kernel size of 3. All SparseConv3D layers adopt a kernel size of 3, stride of 2, and padding of 1 to downsample the feature map resolution. The channel configurations within VDM are set as $(64,32)$, $(32,64)$, and $(64,128)$.
For LION and DSVT, we follow the default parameter settings from their original papers without modification. Additionally, we conduct an only-diffusion experiment, in which the voxel size along the $x$ and $y$ axes is scaled by a factor of 4, while the $z$ axis remains unchanged.
Training Process. We train all models using 8 NVIDIA A40 GPUs. On the Waymo Open Dataset, the VDM-SSMs configuration is trained for 24 epochs with a batch size of 16, while the VDM-DSVT configuration uses 24 epochs with a batch size of 8. For the remaining three datasets, we adopt the same training strategy as that used in the LION (Liu et al. 2024) baseline.
Main Results
To verify the effectiveness of the proposed module, we design two models: VDM-SSMs based on LION and VDM-Transformer based on DSVT. We evaluate their performance on four benchmark datasets: Waymo, nuScenes, Argoverse V2, and ONCE.
Results on WOD. As shown in Table 2, we incorporate the VDM module into both Mamba-based and Transformer-based models. Among them, VDM-Mamba achieves a state-of-the-art (SOTA) result of 74.8 mAPH (L2), outperforming the corresponding baseline LION model (without VDM) by +0.7 mAPH, and also surpassing the recent UniMamba (Jin et al. 2025).
In addition, VDM-DSVT also demonstrates performance gains, achieving a +1.2 mAPH (L2) improvement over the original DSVT, which validates the general applicability of the VDM module to both SSM-based and Transformer-based detectors.
When only the diffusion mechanism is retained in VDM (i.e., VDM-OD-Mamba), the model still achieves the best results on the Vehicle and Cyclist categories, highlighting the importance of the diffusion component in the overall design. Moreover, VDM-Mamba achieves the best performance in the Pedestrian category, suggesting that the fine-grained aggregation ability introduced by VDM is particularly beneficial for small object detection.
Method Vehicle Pedestrian Cyclist mAP
overall 0-30m 30-50m 50m+ overall 0-30m 30-50m 50m+ overall 0-30m 30-50m 50m+
PointRCNN 52.1 74.5 40.9 16.8 4.3 6.2 2.4 0.9 29.8 46.0 20.9 5.5 28.7
PointPillars 68.6 80.9 62.1 47.0 17.6 19.7 15.2 10.2 46.8 58.3 40.3 25.9 44.3
SECOND 71.2 84.0 63.0 47.3 26.4 29.3 24.1 18.1 58.0 70.0 52.4 34.6 51.9
PV-RCNN 77.8 89.4 72.6 58.6 23.5 25.6 22.8 17.3 59.4 71.7 52.6 36.2 53.6
CenterPoint 66.8 80.1 59.6 43.4 49.9 56.2 42.6 26.3 63.5 74.3 57.9 41.5 60.1
PointPainting 66.2 80.3 59.8 42.3 44.8 52.6 36.6 22.5 62.3 73.6 57.2 40.4 57.8
LION 78.2 89.1 72.6 57.5 53.2 62.4 44.0 24.5 68.5 79.2 63.2 47.1 66.6
VDM-mamba 78.6 89.1 74.3 59.3 54.2 63.3 43.2 26.1 69.9 79.4 64.9 49.8 67.6
Table 5: Comparison with previous methods on ONCE validation set.
Results on nuScenes. To further verify the effectiveness of the proposed VDM module, we conduct experiments on both the nuScenes validation and test sets. As shown in Table 3, VDM-mamba achieves 68.1 mAP and 72.5 NDS on the validation set, and 70.0 mAP and 73.7 NDS on the test set.
Although VDM-mamba slightly underperforms UniMamba in terms of overall detection accuracy, it achieves the best results in the pedestrian and bicycle categories on the nuScenes validation set. These categories are particularly important for autonomous driving applications, demonstrating that our proposed method offers clear advantages in detecting small and vulnerable objects. Furthermore, VDM-OD-Mamba achieves a state-of-the-art 72.9 NDS, suggesting that the diffusion-only structure plays a critical role in improving detection performance.
Results on Argoverse V2. As shown in Table 4, our proposed models, VDM-Mamba and VDM-OD-Mamba, achieve state-of-the-art detection performance on the Argoverse 2 validation set, reaching 42.3 mAP and 42.6 mAP, respectively. Compared with the strong baseline LION (Liu et al. 2024), VDM-Mamba yields a notable improvement of 0.8 mAP. Notably, VDM-OD-Mamba, which retains only the diffusion structure within the VDM module, still surpasses all previous approaches, highlighting the effectiveness of the diffusion mechanism. These results further demonstrate the strong representational capacity and scalability of the VDM module in large-scale point cloud scenarios.
Results on ONCE. VDM-Mamba also demonstrates superior performance on the ONCE dataset. It achieves detection accuracies of 78.6 mAP, 54.2 mAP, and 69.9 mAP on the Vehicle, Pedestrian, and Cyclist categories, respectively. The overall detection accuracy across all categories reaches 67.6 mAP, representing a +1.0 mAP improvement over the Mamba-based LION detector. These results establish VDM-Mamba as the new state-of-the-art (SOTA) on the ONCE benchmark.
Ablation Study
Diffusion Waymo(3D AP/APH)L2
Vehicle Pedestrian Cyclist
No(LION) 68.8/68.3 77.3/72.6 74.7/73.8
Yes(VDM-OD) 70.3/69.9 78.2/73.8 76.0/75.0
Table 6: Detection results of the voxel diffusion module on the Waymo validation set (20% data). ”No” indicates that the original voxels are not diffused, and the detection results are entirely determined by LION(Liu et al. 2024) ”Yes” indicates that we only adopt the diffusion module from VDM, i.e., using sparse 3D spconv with a kernel size of 3 and a stride of 1.
Ablation Study on Voxel Diffusion. To assess the contribution of voxel diffusion to detection performance, we conduct ablation experiments on the Waymo validation set using a fixed voxel size of [0.32, 0.32, 0.1875]. We evaluate the model both with and without voxel diffusion. As shown in Table 6, introducing voxel diffusion improves the VDM-OD model by 1.3 mAPH (L2), underscoring its critical role in sequential detection frameworks. Moreover, on the nuScenes (Table 3) and Argoverse 2 (Table 4) datasets, we observe that under the Only-diffusion setting, the model still achieves higher detection accuracy, further validating the significant impact of voxel diffusion.
Fine-grained ONCE (Overall) Waymo (Cyclist)
Vehicle Pedestrian Cyclist mAP AP (L1) APH (L1) AP (L2) APH (L2)
Base (VDM-OD) 78.5 51.0 68.7 66.1 78.5 77.6 75.7 74.7
Fine (VDM) 78.6 54.2 69.9 67.6 79.1 78.1 76.2 75.2
Table 7: Ablation study on the ONCE and Waymo (20% data) dataset for fine-grained feature aggregation. Base denotes the detection results obtained by combining the VDM and Mamba network without fine voxel partitioning. Fine represents the results after applying fine-grained voxel partitioning.
Ablation on Fine-grained Feature Aggregation. To further evaluate the fine-grained feature aggregation capability of VDM under more detailed voxel partitioning, we conducted experiments by refining the voxel resolution. Specifically, we reduced the voxel size along the x and y axes to one-fourth of the original size, while keeping the resolution along the z-axis unchanged, with voxel diffusion still enabled.
As shown in Table 2, under this fine-grained aggregation setting, the model achieves a 0.5 APH (L2) improvement in detecting the Pedestrian category, indicating enhanced localization and recognition of small-scale objects. Furthermore, as reported in Table 7, on the ONCE dataset, the VDM-OD model exhibits inferior performance across all three categories when only voxel diffusion is applied. With the integration of fine-grained feature aggregation, the model achieves a 1.5 mAP gain, demonstrating the effectiveness of finer voxel partitioning in enhancing detection performance. When trained with 20% of the Waymo validation set, the VDM model also achieves higher detection accuracy for the Cyclist category.
Conclusion
We propose the Voxel Diffusion Module (VDM), which consists of a sparse 3D convolutional network, sparse residual blocks, and submanifold sparse convolution (SubM3D). This module diffuses voxelized point clouds via sparse 3D convolutions to increase the density of foreground voxels for more accurate object localization, while simultaneously performing spatial pre-aggregation of voxel features to provide spatial context for the subsequent serialized voxel sequences. To validate the effectiveness of our approach, we evaluate two model variants: VDM-LION (based on Mamba) and VDM-DSVT (based on Transformer). Experimental results demonstrate that VDM consistently improves performance in both SSMs-based and Transformer-based architectures. Notably, VDM-LION (SSMs) achieves state-of-the-art results on the Waymo, Argoverse 2, and ONCE datasets.

### v3

@@
-A Unified Voxel Diffusion Module for Point Cloud 3D Object Detection
+Voxel Densification for Serialized 3D Object Detection: Mitigating Sparsity via Pre-serialization Expansion
 Abstract
-Recent advances in point cloud object detection have increasingly adopted Transformer-based and State Space Models (SSMs), demonstrating strong performance.
-However, voxel-based representations in these models require strict consistency in input and output dimensions due to their serialized processing, which limits the spatial diffusion capability typically offered by convolutional operations.
-This limitation significantly affects detection accuracy.
-Inspired by CNN-based object detection architectures, we propose a novel Voxel Diffusion Module (VDM) to enhance voxel-level representation and diffusion in point cloud data.
-VDM is composed of sparse 3D convolutions, submanifold sparse convolutions, and residual connections.
-To ensure computational efficiency, the output feature maps are downsampled to one-fourth of the original input resolution.
-VDM serves two primary functions: (1) diffusing foreground voxel features through sparse 3D convolutions to enrich spatial context, and (2) aggregating fine-grained spatial information to strengthen voxel-wise feature representation.
-The enhanced voxel features produced by VDM can be seamlessly integrated into mainstream Transformer- or SSM-based detection models for accurate object classification and localization, highlighting the generalizability of our method.
-We evaluate VDM on several benchmark datasets by embedding it into both Transformer-based and SSM-based models.
-Experimental results show that our approach consistently improves detection accuracy over baseline models.
-Specifically, VDM-SSMs achieve 74.7 mAPH (L2) on Waymo, 72.9 NDS on nuScenes, 42.3 mAP on Argoverse 2, and 67.6 mAP on ONCE, setting new state-of-the-art performance across all datasets.
-Our code will be made publicly available.
-Introduction
-3D object detection based on point clouds plays a crucial role in applications such as autonomous driving (Bansal, Krizhevsky, and Ogale 2018) and robotic navigation (Zhu et al.
-2017).
-Inspired by the remarkable success of Transformers (Vaswani et al.
-2017) and State Space Models (SSMs) (Gu and Dao 2023) in natural language processing, researchers have recently explored adapting these architectures to computer vision tasks, achieving strong performance.
-For example, in the domain of point cloud 3D object detection, DVST (Wang et al.
-2023) introduces a dynamic sparse window attention mechanism that adaptively partitions each window into local sub-regions based on voxel sparsity, enabling fully parallel attention computations.
-LION (Liu et al.
-2024) further extends this direction by employing large grouping windows and leveraging SSMs’ ability to capture long-range voxel dependencies, setting new state-of-the-art (SOTA) performance on several public datasets.
-Figure 1: Performance comparison of different models on various datasets. Numbers highlighted in red indicate that our model achieves state-of-the-art results.
-Despite these advancements, both Transformer-based and SSM-based models for point cloud voxel sequence processing require consistency in input and output sequence dimensions, which fundamentally limits their ability to perform convolution-like spatial diffusion.
-However, prior work (Zhang et al.
-2024a; Liu et al.
-2024) has demonstrated that voxel diffusion, particularly for foreground voxels, is critical for enhancing detection accuracy.
-For instance, LION selects top-$k$ voxels via feature summation and applies directional diffusion during intermediate processing.
-In contrast, our method proposes to conduct voxel diffusion before serialization, using convolutional operations to enrich the original voxel space with contextually relevant features.
-Motivated by CNN-based object detection architectures, we propose a novel Voxel Diffusion Module (VDM) tailored for both Mamba-based and Transformer-based 3D detection models.
-Positioned before voxel sequence serialization, VDM leverages 3D convolutional operations to diffuse foreground voxel features and increase their spatial density.
-Simultaneously, sparse convolutions are employed to aggregate fine-grained spatial features within each voxel’s local 3D neighborhood, supplying the downstream sequence model with richer spatial context.
-Structurally, VDM consists of stacked sparse 3D convolutions and residual blocks, effectively supporting both diffusion and aggregation.
-The output feature map is downsampled to one-fourth of the input resolution (in height and width), helping reduce the computational load for subsequent models.
+Recent advances in point cloud object detection have increasingly adopted Transformer-based and State Space Models (SSMs) to capture long-range dependencies.
+However, these serialized frameworks strictly maintain the consistency of input and output voxel dimensions, inherently lacking the capability for voxel expansion.
+This limitation hinders performance, as expanding the voxel set is known to significantly enhance detection accuracy, particularly for sparse foreground objects.
+To bridge this gap, we propose a novel Voxel Densification Module (VDM).
+Unlike standard convolutional stems, VDM is explicitly designed to promote pre-serialization spatial expansion.
+It leverages sparse 3D convolutions to propagate foreground semantics to neighboring empty voxels, effectively densifying the feature representation before it is flattened into a sequence.
+Simultaneously, VDM incorporates residual sparse blocks to aggregate fine-grained local context, ensuring rich geometric feature extraction.
+To balance the computational overhead of increased voxel density, we introduce a strategic cascaded downsampling mechanism.
+We integrate VDM into both Transformer-based (DSVT) and SSM-based (LION) detectors.
+Extensive experiments demonstrate that VDM consistently improves detection accuracy across multiple benchmarks.
+Specifically, our method achieves 74.8 mAPH (L2) on the Waymo validation set and 70.5 mAP on the nuScenes test set.
+Furthermore, it attains 42.6 mAP on the Argoverse 2 validation set and 67.6 mAP on the ONCE validation set, consistently outperforming the baseline models.
+The source code will be made publicly available at https://github.com/qifeng22/VDM.
+I Introduction
+3D object detection based on point clouds plays a crucial role in applications such as autonomous driving [2] and robotic navigation [34].
+Inspired by the remarkable success of Transformers [22] and State Space Models (SSMs) [8] in natural language processing, researchers have recently explored adapting these architectures to computer vision tasks, achieving strong performance.
+For example, in the domain of point cloud 3D object detection, DSVT [24] introduces a dynamic sparse window attention mechanism that adaptively partitions each window into local sub-regions based on voxel sparsity, enabling fully parallel attention computations.
+LION [14] further extends this direction by employing large grouping windows and leveraging SSMs’ ability to capture long-range voxel dependencies, achieving impressive performance on several public datasets.
+Despite these advancements, a fundamental limitation persists in current serialized frameworks: the spatial-serial gap.
+Both Transformer-based and SSM-based models typically flatten sparse voxels into 1D sequences to model long-range dependencies.
+However, this serialization process strictly maintains the input-output token consistency, meaning the set of active voxels remains static throughout the backbone.
+Consequently, these models lack the inherent capability to perform spatial expansion that is trivial in convolutional networks.
+This limitation is critical because raw point clouds are often sparse and incomplete; without expanding the voxel set to include neighboring empty regions, the model struggles to capture sufficient local context for accurate localization, particularly for distant or occluded objects.
+While prior works like LION [14] attempt to mitigate this via intermediate feature diffusion, they operate on already-serialized features, where spatial topology is disrupted.
+In contrast, we argue that the optimal strategy is to explicitly densify the voxel representation before serialization, thereby bridging the gap between spatial completeness and sequential modeling efficiency.
+Figure 1: Comparison of serialized detection paradigms.
+(a) Existing methods such as LION [14] and DSVT [24] directly process sparse voxel sets.
+(b) Our proposed framework incorporates the Voxel Densification Module (VDM) to explicitly densify and aggregate voxel features before sequence modeling.
+Motivated by the spatial flexibility of CNN-based architectures, we propose a novel Voxel Densification Module (VDM), as illustrated in Figure 1, specifically tailored to serve as a pre-serialization stem for Mamba-based and Transformer-based 3D detectors.
+Positioned prior to the flattening of voxel sequences, VDM leverages sparse 3D convolutional operations to explicitly expand the foreground voxel set, effectively densifying the feature representation.
+This process generates new informative voxels in the 3D neighborhood of the original sparse points, ensuring that the subsequent serialized model receives a spatially enriched input.
+Structurally, VDM integrates stacked sparse 3D convolutions with residual blocks to achieve two goals simultaneously: (1) Voxel Densification, which propagates semantics to empty spatial neighbors, and (2) Fine-grained Aggregation, which captures local geometric details.
+Crucially, to balance the computational cost of processing increased voxel counts, we introduce a strategic downsampling mechanism (reducing resolution to 1/4), optimizing the trade-off between spatial density and sequence length.
 VDM is designed to be modular and general, allowing seamless integration with both Transformer-based and SSM-based detectors.
 To evaluate its effectiveness, we perform extensive experiments on four large-scale datasets.
-As shown in Figure 1, our VDM-enhanced SSM-based model (VDM-Mamba) achieves 74.7 mAPH (L2) on Waymo, 68.1 mAP on nuScenes, 42.3 mAP on Argoverse 2, and 67.6 mAP on ONCE, setting a new state-of-the-art (SOTA) performance.
-Additionally, the Transformer-based variant (VDM-DSVT) improves performance by +2.2 mAPH (L2) over the DSVT baseline on the Waymo dataset, confirming the general effectiveness of our approach.
+Our VDM-enhanced SSM-based frameworks achieve 74.8 mAPH (L2) on the Waymo validation set, 70.5 mAP on the nuScenes test set, 42.6 mAP on the Argoverse 2 validation set, and 67.6 mAP on the ONCE validation set, consistently outperforming baseline models across different benchmarks.
+Additionally, the Transformer-based variant (VDM-DSVT) improves performance by +1.2 mAPH (L2) over the DSVT baseline on the Waymo dataset, confirming the general effectiveness of our approach.
 Our contributions can be summarized as follows:
-  • We propose VDM, a voxel diffusion module that enhances the representation of raw voxel inputs through 3D convolution, significantly enriching foreground voxel features and improving detection accuracy.
-  • VDM utilizes sparse 3D convolutions and residual networks to aggregate spatial features from the voxel neighborhood, providing richer positional cues for downstream sequence modeling.
-  • VDM is a general-purpose module that integrates well with both Transformer-based and SSM-based 3D detectors, consistently boosting model performance.
-  • We validate the effectiveness of VDM through experiments on multiple public datasets, where VDM-SSMs achieve SOTA results, and VDM also improves Transformer-based models.
-Related Work
-3D object detection based on point clouds has seen significant progress in recent years.
-According to model architecture, existing approaches can be broadly categorized into three types: CNN-based, Transformer-based, and SSM-based methods.
-CNN-based models emerged earlier and have been thoroughly studied, achieving solid performance.
-In contrast, Transformer and Mamba-based models are relatively recent, and while less explored for 3D object detection, they have already shown promising results—often surpassing CNN-based methods in detection accuracy.
+  • We identify the limitation of static voxel sets in serialized 3D detectors and propose VDM (Voxel Densification Module), a plug-and-play operator that explicitly densifies sparse voxel inputs via pre-serialization convolution.
+  • We design a dual-path architecture within VDM that simultaneously performs spatial expansion (to increase foreground recall) and fine-grained feature aggregation (to enhance local context).
+  • VDM demonstrates universal compatibility, seamlessly integrating with both Transformer-based and SSM-based backbones to significantly boost performance.
+  • We validate the effectiveness of VDM through extensive experiments, achieving state-of-the-art results across Waymo, nuScenes, Argoverse 2, and ONCE datasets.
+II Related Work
+3D object detection based on point clouds has seen significant progress.
+Existing approaches can be broadly categorized into CNN-based, Transformer-based, and State Space Model (SSM)-based methods.
+While CNN-based models have established a strong baseline, recent serialization-based architectures (Transformers and SSMs) have shown superior potential in capturing long-range dependencies, though they face distinct challenges regarding spatial continuity.
 CNN-Based 3D Object Detection.
-CNN-based 3D detectors are typically categorized into voxel-based and point-based approaches.
-VoxelNet (Zhou and Tuzel 2018) introduces a Voxel Feature Encoding (VFE) layer to convert point clusters within voxels into unified feature representations.
-HEDNet uses an encoder-decoder design to capture long-range spatial dependencies, particularly useful for detecting large or distant objects.
-SAFDNet proposes an adaptive feature diffusion strategy to address missing center features in sparse detectors.
-PointPillars (Lang et al.
-2019) leverages PointNet to learn features from vertical columns of point clouds, offering notable inference speed advantages.
-CenterPoint (Yin, Zhou, and Krahenbuhl 2021) detects object centers using a keypoint detector and regresses attributes such as 3D size, orientation, and velocity.
-3DSSD (Yang et al.
-2020) introduces a fused sampling strategy in the downsampling process to retain informative representative points.
-IASSD (Zhang et al.
-2022) proposes two task-aware, instance-sensitive sampling strategies for hierarchically selecting foreground points of interest.
-Among these approaches, voxel-based detectors generally achieve higher accuracy and have become the dominant choice for outdoor 3D detection tasks.
+CNN-based detectors are the pioneers in this field, utilizing either voxel-based or point-based representations.
+VoxelNet [32] introduces Voxel Feature Encoding (VFE) to generate unified feature representations.
+To improve efficiency, PointPillars [12] collapses point clouds into vertical columns, enabling fast 2D convolution processing.
+CenterPoint [27] advances this by adopting a center-based anchor-free detection head.
+More recently, methods like HEDNet [29] utilize encoder-decoder designs for long-range perception.
+Notably, SAFDNet [28] identifies the issue of feature sparsity and proposes an adaptive feature diffusion strategy.
+However, SAFDNet primarily focuses on diffusing features within the backbone for CNN-based detectors.
+In contrast, our work targets the specific ”spatial-serial gap” in serialization-based models, aiming to physically densify the voxel set prior to sequence flattening.
 Transformer-Based 3D Object Detection.
-Transformers have achieved state-of-the-art performance in both NLP and computer vision tasks.
-PVT-SSD (Yang et al.
-2023) introduces a Point-Voxel Transformer that captures long-range dependencies via voxel tokens while preserving fine spatial detail through point-level features.
-FlatFormer (Liu et al.
-2023) flattens point clouds using window-based sorting and partitions them into equal-sized groups, improving detection efficiency.
-DSVT (Wang et al.
-2023) proposes a Dynamic Sparse Window Attention mechanism that adaptively splits each window based on voxel sparsity, enabling fully parallel feature computation.
-However, the serialization of voxel features in Transformer-based methods can lead to the loss of fine-grained spatial cues, limiting performance in complex 3D scenes.
+Transformers have demonstrated state-of-the-art performance by treating voxels as sequences.
+PVT-SSD [26] employs a Point-Voxel Transformer to capture global dependencies.
+FlatFormer [15] improves efficiency by flattening point clouds via window-based sorting.
+DSVT [24] introduces Dynamic Sparse Window Attention to handle sparsity efficiently.
+Despite their success, these methods typically operate on a static set of non-empty voxels.
+Due to the strict input-output consistency of attention mechanisms, they lack the inherent ability to expand the voxel set to neighboring empty regions, often leading to insufficient context for sparse objects.
 SSM-Based 3D Object Detection.
-SSM-based detectors represent a new and emerging line of work.
-LION (Liu et al.
-2024) applies a large-group strategy during voxel serialization for more effective spatial context modeling.
-Voxel Mamba (Zhang et al.
-2024b) eliminates the grouping step and serializes the entire voxel space as a single sequence, introducing a Dual-Scale SSM Block to capture hierarchical spatial dependencies.
-UniMamba combines 3D convolution with SSMs in a compact multi-head architecture that efficiently captures both local and global spatial context.
-Although still in their early stages, SSM-based models already outperform traditional CNN-based frameworks on several benchmarks due to their strong modeling capabilities.
-Our proposed Voxel Diffusion Module (VDM) further enhances these serialized architectures by introducing voxel diffusion and fine-grained spatial feature aggregation prior to serialization.
-This design supplies Transformer- and SSM-based models with richer input representations, leading to significant performance gains in 3D object detection.
+Mamba-based SSMs represent a burgeoning direction.
+LION [14] leverages SSMs with large window grouping to model wider spatial contexts.
+Voxel Mamba [30] serializes the entire voxel space to capture global dependencies via Dual-Scale SSM Blocks.
+UniMamba [11] integrates 3D convolutions with SSMs in a multi-head architecture.
+While UniMamba combines CNNs and SSMs, it treats them as parallel or interleaved feature extractors.
+Different from these approaches, our proposed Voxel Densification Module (VDM) is explicitly designed as a pre-serialization stem.
+Instead of merely extracting features, VDM fundamentally alters the spatial distribution of the input by densifying foreground voxels.
+This ensures that the downstream serialized models (whether Transformer or SSM) receive a spatially enriched token sequence, directly addressing the sparsity limitation inherent in serialization.
 Figure 2: Overall architecture of our proposed method.
-(a) The Voxel Diffusion Module (VDM) is placed before the sequential detection model, enabling diffusion of raw voxelized point clouds and aggregation of spatial voxel features.
-Internally, it consists of Submanifold 3D Convolution (SubM3D), Sparse Residual Block (SRB-3D), and Sparse 3D Convolution (SPConv-3D), where SPConv-3D plays a key role in voxel diffusion.
-(b) Mamba Block.
-(c) Sparse Residual Block (SRB-3D).
-The red connection line ”Only-Diffusion” in the figure indicates the model architecture when only the voxel diffusion strategy is applied in the VDM, aiming to highlight the critical impact of voxel diffusion on detection performance.
-Method
-In this section, we introduce the proposed Voxel Diffusion Module (VDM), designed to enhance the representation of voxelized point cloud features.
-VDM leverages convolutional operations to perform early-stage voxel diffusion, increasing the density of meaningful voxel information.
+(a) The Voxel Densification Module (VDM) is positioned prior to serialization, specifically designed to densify sparse voxel inputs and aggregate local context.
+The core VDM path (gray box) utilizes stacked Submanifold 3D Convolutions (SubM-3D) and Sparse Residual Blocks (SRB-3D) with downsampling layers (Spconv-3D) to reduce resolution to 1/4.
+(b) The Mamba Block used in SSM-based backbones.
+(c) The detailed structure of the Sparse Residual Block (SRB-3D).
+Note: The red path labeled “Only-Densification” represents the ablation variant where only a single expansion layer (Spconv-3D with $K=3,S=1$) is applied, isolating the efficacy of the voxel expansion mechanism.
+III Method
+In this section, we introduce the proposed Voxel Densification Module (VDM), designed to enhance the representation of voxelized point cloud features.
+VDM leverages convolutional operations to perform early-stage voxel expansion, increasing the density of meaningful voxel information.
 This provides richer input for downstream serialized models such as Transformers or Mamba.
 Furthermore, VDM aggregates local spatial features within each voxel’s 3D neighborhood, enabling the sequence models to more effectively utilize spatial positional cues during processing.
-Preliminaries
+III-A Preliminaries
 Transformer. Transformers, based on self-attention mechanisms, have been widely adopted in both natural language processing and computer vision. The standard attention mechanism is defined as:
@@
 Here, $Q$ (Query), $K$ (Key), and $V$ (Value) represent the input projection vectors.
-$Q$ interacts with $K^{\top}$ to compute similarity scores, which are then used to weight the values in $V$.
-The term $d_{k}$ denotes the dimensionality of the key vectors, and the scaling factor $\sqrt{d_{k}}$ is introduced to stabilize the gradients during training.
 While Transformers are powerful in capturing global dependencies, their quadratic time complexity poses a challenge when handling long sequences, such as in voxelized point clouds.
@@
 where $A^{\prime}$ and $B^{\prime}$ are obtained from continuous $A$ and $B$ via a discretization rule: $A^{\prime}=f_{A}(\Delta,A)$ and $B^{\prime}=f_{B}(\Delta,A,B)$.
-Mamba-based SSM models (Gu and Dao 2023) have demonstrated significant advantages over Transformers in terms of modeling long-range dependencies and computational efficiency.
-In the context of 3D point cloud detection, such models (Jin et al.
-2025; Liu et al.
-2024) have shown promising results and even outperformed many Transformer-based architectures.
+Mamba-based SSM models [8] have demonstrated significant advantages over Transformers in terms of modeling long-range dependencies and computational efficiency.
+In the context of 3D point cloud detection, such models [11, 14] have shown promising results and even outperformed many Transformer-based architectures.
 Sparse Residual Block (SRB).
-In CNN-based 3D detectors such as SAFDNet (Zhang et al.
-2024a), the Sparse Residual Block (SRB) is commonly used for efficient feature extraction.
-An SRB consists of two 3D submanifold sparse convolution layers and a residual skip connection linking the input to the output.
-In our design, VDM is composed of stacked SRB modules and additional SubM3D layers to enable both voxel diffusion and spatial feature aggregation.
-This structure facilitates the enrichment of voxel-level spatial representations before they are passed into the serialized backbone (Transformer or SSM), thereby improving detection accuracy.
-Figure 3: Visualization of the Voxel Diffusion Module (VDM) applied to the original point clouds, demonstrating its effect in voxel-level diffusion and feature aggregation.
-Overall Architecture
-As illustrated in Figure 2, the Voxel Diffusion Module (VDM) is designed to perform voxel-level diffusion of point clouds and aggregate fine-grained voxel features.
-This facilitates the generation of rich foreground voxel representations and fine-grained spatial features, serving as a beneficial foundation for subsequent transformer-based and state-space model (SSM)-based detectors.
-Convolutional neural network (CNN)-based models (Zhang et al.
-2024a) have been extensively studied in 3D object detection.
-These models typically begin by applying feature extraction networks to voxelized point clouds, followed by a backbone and a detection head for final predictions.
-Inspired by this paradigm, the proposed VDM is composed of alternating Submanifold 3D Convolution (SubM3D) layers and Sparse Residual Blocks (SRB).
-Additionally, two Sparse 3D Convolution (SPConv3D) layers are inserted to downsample and upsample the voxel feature maps, enabling hierarchical feature learning and effective diffusion.
-Voxel Diffusion in VDM.
-Voxel diffusion plays a significant role in enhancing model detection performance, as demonstrated in previous studies (Liu et al.
-2024; Zhang et al.
-2024a).
-Inspired by the property of convolution operations to increase the number of sparse features, we introduce a sparse 3D convolution with a kernel size of 3 within our proposed Voxel Diffusion Module (VDM).
-This module enables diffusion over the input voxel grid, thus expanding the spatial distribution of informative voxels.
-Figure 3 illustrates how VDM diffuses the original point cloud into a denser voxel representation, providing a richer feature basis for subsequent serial detection networks.
-To quantitatively evaluate this effect, we perform a statistical analysis on the voxel count post-VDM processing, as shown in Table 1.
+In CNN-based 3D detectors such as SAFDNet [28], the Sparse Residual Block (SRB) is commonly used for efficient feature extraction.
+In our design, VDM is composed of stacked SRB modules and additional SubM3D layers to enable both voxel expansion and spatial feature aggregation.
+This structure facilitates the enrichment of voxel-level spatial representations before they are passed into the serialized backbone.
+III-B Overall Architecture
+As illustrated in Figure 2, the Voxel Densification Module (VDM) is designed to perform voxel-level densification of point clouds and aggregate fine-grained voxel features.
+This facilitates the generation of rich foreground voxel representations, serving as a beneficial foundation for subsequent Transformer-based and SSM-based detectors.
+Inspired by CNN-based paradigms, the proposed VDM is composed of alternating Submanifold 3D Convolution (SubM3D) layers and Sparse Residual Blocks (SRB).
+Additionally, two Sparse 3D Convolution (SPConv3D) layers are inserted to downsample the voxel feature maps, enabling hierarchical feature learning and effective spatial expansion.
+Figure 3: Visualization of the Voxel Densification Module (VDM) applied to the original point clouds, demonstrating its effect in voxel-level expansion and feature aggregation.
+Voxel Densification in VDM.
+Voxel densification plays a significant role in enhancing model detection performance.
+Inspired by the property of convolution operations to increase the number of sparse features (dilation), we introduce a sparse 3D convolution with a kernel size of 3 within our proposed VDM.
+This module enables expansion over the input voxel grid, thus enlarging the spatial distribution of informative voxels.
+Figure 3 illustrates how VDM expands the original point cloud into a denser voxel representation, providing a richer feature basis for subsequent serial detection networks.
+To quantitatively evaluate this effect, we perform a statistical analysis on the voxel count post-VDM processing, as shown in Table I.
 As observed, the number of total and foreground voxels significantly increases after applying VDM.
-This expansion ensures that subsequent modules—particularly those based on transformers or State Space Models (SSMs)—can access more comprehensive foreground point features, thereby improving object detection accuracy.
-It is worth noting that voxel diffusion serves as a crucial component of the VDM.
-Our later experiments will further highlight its contribution to performance gains, underlining its importance in boosting the effectiveness of serialization-based architectures such as transformers and Mamba variants.
+This expansion ensures that subsequent modules—particularly those based on Transformers or SSMs—can access more comprehensive foreground point features, thereby improving object detection accuracy.
+It is worth noting that voxel expansion serves as a crucial component of the VDM.
+Our later experiments will further highlight its contribution to performance gains.
 Figure 4: Fine-grained feature aggregation within the VDM, utilizing convolutional operations to aggregate relative spatial information across voxels.
-Method Overall Voxel Count Foreground Voxel Count mAPH L2
-Non(LION) 36340 2616 74.0
-VDM 188562 9146 74.8
-VDM(OD) 216601 10705 74.8
-Table 1: Effect of VDM on voxel counts in input point clouds on the Waymo validation dataset.
-The overall and foreground voxel counts are computed as the total number of corresponding voxels across all frames divided by the total number of frames.
-Methods Operator Vehicle 3D AP/APH Pedestrian 3D AP/APH Cyclist 3D AP/APH mAP/mAPH (L2)
-L1 L2 L1 L2 L1 L2
-SECOND
-CNN
-72.3/71.7 63.9/63.3 68.7/58.2 60.7/51.3 60.6/59.3 58.3/57.0 61.0/57.2
-PointPillars 72.1/71.5 63.6/63.1 70.6/56.7 62.8/50.3 64.4/62.3 61.9/59.9 62.8/57.8
-CenterPoint 74.2/73.6 66.2/65.7 76.6/70.5 68.8/63.2 72.3/71.1 69.7/68.5 68.2/65.8
-PV-RCNN 78.0/77.5 69.4/69.0 79.2/73.0 70.4/64.7 71.5/70.3 69.0/67.8 69.6/67.2
-PillarNet-18 78.2/77.7 70.4/69.9 79.8/72.6 71.6/64.9 70.4/69.3 67.8/66.7 69.9/67.2
-FSD 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3 72.9/70.8
-AFDetV2 77.6/77.1 69.7/69.2 80.2/74.6 72.2/67.0 73.7/72.7 71.0/70.1 71.0/68.8
-PillarNeXt 78.4/77.9 70.3/69.8 82.5/77.1 74.9/69.8 73.2/72.2 70.6/69.6 71.9/69.7
-VoxelNext 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2 72.2/70.1
-CenterFormer 75.0/74.4 69.9/69.4 78.6/73.0 73.6/68.3 72.3/71.3 69.8/68.8 71.1/68.9
-PV-RCNN++ 79.3/78.8 70.6/70.2 81.3/76.3 73.2/68.0 73.7/72.7 71.2/70.2 71.7/69.5
-ConQueR 76.1/75.6 68.7/68.2 79.0/72.3 70.9/64.7 73.9/72.5 71.4/70.1 70.3/67.7
-FocalFormer3D –/– 68.1/67.6 –/– 72.7/66.8 –/– 73.7/72.6 71.5/69.0
-HEDNet 81.1/80.6 73.2/72.7 84.4/80.0 76.8/72.6 78.7/77.7 75.8/74.9 75.3/73.4
-VPF 80.2/79.7 71.9/71.5 82.5/76.9 74.8/69.4 77.1/76.0 74.2/73.2 73.6/71.4
-TransFusion
-Transformer
-–/– –/65.1 –/– –/63.7 –/– –/65.9 –/64.9
-SST_TS 76.2/75.8 68.0/67.6 81.4/74.0 72.8/65.9 –/– –/– –/–
-SWFormer 77.8/77.3 69.2/68.8 80.9/72.7 72.5/64.9 –/– –/– –/–
-FlatFormer - 69.0/68.6 - 71.5/65.3 –/– 68.6/67.5 69.7/67.2
-OcTr 78.1/77.6 69.8/69.3 80.8/74.4 72.5/66.5 72.6/71.5 69.9/68.9 70.7/68.2
-DSVT-Pillar 79.3/78.8 70.9/70.5 82.8/77.0 75.2/69.8 76.4/75.4 73.6/72.7 73.2/71.0
-DSVT-Voxel 79.2/78.8 70.9/70.5 82.9/77.7 75.2/70.3 77.2/76.2 74.3/73.4 73.5/71.4
-VDM-DSVT(Ours) 79.7/79.3 71.4/71.0 84.3/79.7 76.7/72.3 78.4/77.4 75.6/74.6 74.6/72.6
-Voxel Mamba
-SSMs
-80.8/80.3 72.6/72.2 85.0/80.8 77.7/73.6 78.6/77.6 75.7/74.8 75.3/73.6
-LION 80.3/79.9 72.0/71.6 85.8/81.4 78.5/74.3 80.1/79.0 77.2/76.2 75.9/74.0
-UniMamba 80.6/80.1 72.3/71.8 86.0/81.3 78.7/74.1 80.3/79.3 77.5/76.5 76.1/74.1
-VDM-Mamba (Ours) 81.0/80.5 72.6/72.2 86.4/82.5 79.3/75.5 80.4/79.4 77.5/76.5 76.5/74.8
-VDM-OD-Mamba (Ours) 81.3/80.9 73.0/72.6 86.1/82.1 78.9/75.0 80.7/79.7 77.8/76.9 76.6/74.8
-Table 2: Comparison with prior methods (CNNs, Transformers, and SSMs) on the Waymo Open Dataset using 100% training data.
-Metrics are reported as mAP/mAPH (%) $\uparrow$ for overall results and AP/APH (%) $\uparrow$ per category.
-All models are trained under the single-frame setting.
-“–” indicates unavailable results.
-VDM-OD-Mamba denotes a variant where the Only-Diffusion structure is used within the VDM framework.
-The best-performing results are highlighted in bold.
 Fine-Grained Feature Aggregation.
-To capture more detailed spatial information, we further divide the voxel grid into finer partitions, as illustrated in Figure 4.
-Leveraging the strong spatial perception capabilities of the Submanifold 3D Convolution (SubM3D) and the Sparse Residual Block (SRB) within the VDM, we extract fine-grained voxel features from these subdivided regions.
-In addition, we adopt a sparse convolution with a stride of 2 to serve dual purposes: first, it reduces the resolution of the feature map to half, which effectively controls computational cost by mitigating the overhead introduced by voxel diffusion; second, it aggregates local fine-grained features, thereby enriching the spatial context available for downstream modules.
-This design ensures that the model benefits from both detailed spatial representations and efficient computation during object detection.
-The function of VDM is to diffuse the number of sparse voxels and to aggregate fine-grained features. It can be summarized by the following formulas:
-$\mathit{FG}^{\prime}=\mathrm{Voxel\_Diffusion}(\mathit{FG},\mathit{spconv})$ (6)
-$\mathit{F}^{\prime}=\mathrm{Aggregation}(\mathit{F},\mathit{subm},\mathit{spconv})$ (7)
-where $\mathit{F}$ denotes the initial voxel feature map; $\mathit{F}^{\prime}$ represents the feature map after local voxel aggregation, with width and height reduced to one-fourth of the original size; $\mathit{subm}$ denotes submanifold sparse convolution, and $\mathit{spconv}$ denotes sparse convolution.
-$\mathit{FG}^{\prime}$ indicates the number of foreground voxels after VDM processing, and $\mathit{FG}$ is the number of foreground voxels in the initial feature map.
+To capture intricate spatial details, we design VDM as a hierarchical framework.
+As shown in Figure 2(a), the input sparse voxel grid is first processed by a Submanifold 3D Convolution (SubM-3D) stem to generate the initial feature embedding.
+Let $\mathbf{V}^{(0)}=(\mathcal{P}^{(0)},\mathcal{F}^{(0)})$ denote this initial state.
+Here, $\mathcal{P}^{(0)}=\{p_{i}\mid p_{i}\in\mathbb{Z}^{3}\}_{i=1}^{N_{0}}$ represents the set of discrete coordinates for $N_{0}$ active voxels, and $\mathcal{F}^{(0)}\in\mathbb{R}^{N_{0}\times C}$ denotes the corresponding feature vectors.
+To achieve effective voxel densification and context aggregation, we subsequently employ a cascaded downsampling strategy consisting of $L=2$ stages.
+Each stage $l$ comprises a sequence of Sparse Residual Blocks (SRB-3D) for feature refinement, followed by a strided Sparse 3D Convolution (SpConv-3D) for spatial expansion.
+Formally, we define the stage transformation $\mathbf{V}^{(l)}=\mathcal{H}_{l}(\mathbf{V}^{(l-1)})$ for $l\in\{1,2\}$.
+The operator $\mathcal{H}_{l}$ is defined as the sequential composition of two mapping functions:
+First, the feature refinement step utilizes the SRB-3D module, denoted as $\Psi(\cdot)$, to extract high-level semantics while maintaining the input topology:
+$\hat{\mathcal{F}}^{(l-1)}=\Psi(\mathcal{F}^{(l-1)})$ (6)
+where $\hat{\mathcal{F}}^{(l-1)}$ denotes the refined features at the current scale.
+Second, the geometric expansion and aggregation are performed simultaneously by the SpConv-3D layer with learnable weights $\mathcal{W}^{(l)}$ and a stride of $s=2$.
+This operation transforms both the geometry and features:
+$\mathcal{P}^{(l)}=\left\{\left\lfloor\frac{p+k}{2}\right\rfloor\;\middle|\;p\in\mathcal{P}^{(l-1)},\,k\in\mathcal{K}\right\}$ (7)
+$\mathcal{F}^{(l)}=\mathcal{W}^{(l)}\ast\hat{\mathcal{F}}^{(l-1)}$ (8)
+Here, $\mathcal{K}$ is the kernel support.
+Notably, despite the reduction in resolution due to the stride, the cardinality of the active voxel set $|\mathcal{P}^{(l)}|$ remains higher than that of a standard baseline at the equivalent voxel size.
+This phenomenon, attributed to the spatial diffusion effect of the kernel $\mathcal{K}$ in Equation 7, explicitly bridges the spatial-serial gap.
+The final output of VDM is the densified sparse tensor at the second stage, i.e., $\mathbf{V}^{\prime}=\mathbf{V}^{(2)}=(\mathcal{P}^{(2)},\mathcal{F}^{(2)})$, which provides the enriched representation (at $1/4$ scale) for serialization.
 Integration with SSMs and Transformer Modules.
@@
 As the sequence length increases, Mamba-based networks are able to adopt larger grouping sizes due to their lower computational complexity compared to Transformer-based counterparts.
-Given the strong performance of LION (Liu et al.
-2024) and DSVT (Wang et al.
-2023) in the field of 3D object detection, we integrate the VDM module into both networks and conduct experiments to validate its effectiveness.
-Experiments
-Datasets and Evaluation Metrics
-Waymo Open Dataset.
-The Waymo Open Dataset (WOD) (Sun et al.
-2020) includes 1,150 scenes, divided into 798 for training, 202 for validation, and 150 for testing.
-Each scene covers a perception range of $150\,\text{m}\times 150\,\text{m}$.
-For evaluation, WOD adopts 3D mean Average Precision (mAP) and mAP weighted by heading accuracy (mAPH).
-Each metric includes two difficulty levels: L1 for objects with more than five points, and L2 for bounding boxes containing one to five points.
-nuScenes Dataset.
-The nuScenes dataset (Caesar et al.
-2020) contains 1,000 scenes, with 750 for training, 150 for validation, and 150 for testing.
-It uses mean Average Precision (mAP) and nuScenes Detection Score (NDS) as evaluation metrics.
-The dataset covers 10 annotated object classes, representing most common categories in traffic scenarios.
-Argoverse 2 Dataset.
-The Argoverse 2 dataset (Chang et al.
-2019) comprises 1,000 sequences, with 700 for training, 150 for validation, and 150 for testing.
-The perception range extends up to 200 meters.
-It uses mean Average Precision (mAP) as the evaluation metric.
-ONCE Dataset.
-The ONCE dataset (Mao et al.
-2021) consists of 5,000, 3,000, and 8,000 frames for training, validation, and testing respectively.
-It defines three detection classes: Vehicle, Pedestrian, and Cyclist.
-The final performance is evaluated using mean Average Precision (mAP) across all three classes.
+Given the strong performance of LION [14] and DSVT [24] in the field of 3D object detection, we integrate the VDM module into both networks and conduct experiments to validate its effectiveness.
+Method Overall Voxel Count Foreground Voxel Count mAPH L2
+Non(LION) 36340 2616 74.0
+VDM 188562 9146 74.8
+VDM-OD 216601 10705 74.8
+TABLE I: Effect of VDM on voxel counts in input point clouds on the Waymo validation dataset.
+The overall and foreground voxel counts are computed as the total number of corresponding voxels across all frames divided by the total number of frames.
+Methods Operator Vehicle 3D AP/APH Pedestrian 3D AP/APH Cyclist 3D AP/APH mAP/mAPH (L2)
+L1 L2 L1 L2 L1 L2
+SECOND [25] CNN 72.3/71.7 63.9/63.3 68.7/58.2 60.7/51.3 60.6/59.3 58.3/57.0 61.0/57.2
+PointPillars [12] 72.1/71.5 63.6/63.1 70.6/56.7 62.8/50.3 64.4/62.3 61.9/59.9 62.8/57.8
+CenterPoint [27] 74.2/73.6 66.2/65.7 76.6/70.5 68.8/63.2 72.3/71.1 69.7/68.5 68.2/65.8
+PillarNet-18 [17] 78.2/77.7 70.4/69.9 79.8/72.6 71.6/64.9 70.4/69.3 67.8/66.7 69.9/67.2
+FSD [6] 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3 72.9/70.8
+AFDetV2 [9] 77.6/77.1 69.7/69.2 80.2/74.6 72.2/67.0 73.7/72.7 71.0/70.1 71.0/68.8
+VoxelNext [5] 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2 72.2/70.1
+CenterFormer [33] 75.0/74.4 69.9/69.4 78.6/73.0 73.6/68.3 72.3/71.3 69.8/68.8 71.1/68.9
+PV-RCNN++ [19] 79.3/78.8 70.6/70.2 81.3/76.3 73.2/68.0 73.7/72.7 71.2/70.2 71.7/69.5
+HEDNet [29] 81.1/80.6 73.2/72.7 84.4/80.0 76.8/72.6 78.7/77.7 75.8/74.9 75.3/73.4
+VPF [10] 80.2/79.7 71.9/71.5 82.5/76.9 74.8/69.4 77.1/76.0 74.2/73.2 73.6/71.4
+TransFusion [1] Transformer –/– –/65.1 –/– –/63.7 –/– –/65.9 –/64.9
+FlatFormer [15] - 69.0/68.6 - 71.5/65.3 –/– 68.6/67.5 69.7/67.2
+OcTr [31] 78.1/77.6 69.8/69.3 80.8/74.4 72.5/66.5 72.6/71.5 69.9/68.9 70.7/68.2
+DSVT-Pillar [24] 79.3/78.8 70.9/70.5 82.8/77.0 75.2/69.8 76.4/75.4 73.6/72.7 73.2/71.0
+DSVT-Voxel [24] 79.2/78.8 70.9/70.5 82.9/77.7 75.2/70.3 77.2/76.2 74.3/73.4 73.5/71.4
+VDM-DSVT(Ours) 79.7/79.3 71.4/71.0 84.3/79.7 76.7/72.3 78.4/77.4 75.6/74.6 74.6/72.6
+Voxel Mamba [30] SSMs 80.8/80.3 72.6/72.2 85.0/80.8 77.7/73.6 78.6/77.6 75.7/74.8 75.3/73.6
+LION [14] 80.3/79.9 72.0/71.6 85.8/81.4 78.5/74.3 80.1/79.0 77.2/76.2 75.9/74.0
+UniMamba [11] 80.6/80.1 72.3/71.8 86.0/81.3 78.7/74.1 80.3/79.3 77.5/76.5 76.1/74.1
+VDM-Mamba (Ours) 81.0/80.5 72.6/72.2 86.4/82.5 79.3/75.5 80.4/79.4 77.5/76.5 76.5/74.8
+VDM-OD (Ours) 81.3/80.9 73.0/72.6 86.1/82.1 78.9/75.0 80.7/79.7 77.8/76.9 76.6/74.8
+TABLE II: Comparison with prior methods (CNNs, Transformers, and SSMs) on the Waymo Open Dataset validation set using 100% training data.
+Metrics are reported as mAP/mAPH (%) $\uparrow$ for overall results and AP/APH (%) $\uparrow$ per category.
+All models are trained under the single-frame setting.
+“–” indicates unavailable results.
+VDM-OD denotes a variant where the Only-Densification structure is used within the VDM framework.
+The best-performing results are highlighted in bold.
 Performances on the validation set
 Method NDS mAP Car Truck Bus T.L. C.V. Ped. M.T. Bike T.C. B.R.
-CenterPoint 66.5 59.2 84.9 57.4 70.7 38.1 16.9 85.1 59.0 42.0 69.8 68.3
-VoxelNeXt 66.7 60.5 83.9 55.5 70.5 38.1 21.1 84.6 62.8 50.0 69.4 69.4
-Uni3DETR 68.5 61.7
-Transfusion-LiDAR 70.1 65.5 86.9 60.8 73.1 43.4 25.2 87.5 72.9 57.3 77.2 70.3
-DSVT 71.1 66.4 87.4 62.6 75.9 42.1 25.3 88.2 74.8 58.7 77.9 71.0
-HEDNet 71.4 66.7 87.7 60.6 77.8 50.7 28.9 87.1 74.3 56.8 76.3 66.9
-SAFDNet 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
-LION 72.1 68.0 87.9 64.9 77.6 44.4 28.5 89.6 75.6 59.4 80.8 71.6
-UniMamba 72.6 68.5 88.7 64.7 79.7 47.9 28.7 89.7 74.6 59.1 79.5 72.3
+CenterPoint [27] 66.5 59.2 84.9 57.4 70.7 38.1 16.9 85.1 59.0 42.0 69.8 68.3
+Transfusion-LiDAR [1] 70.1 65.5 86.9 60.8 73.1 43.4 25.2 87.5 72.9 57.3 77.2 70.3
+DSVT [24] 71.1 66.4 87.4 62.6 75.9 42.1 25.3 88.2 74.8 58.7 77.9 71.0
+HEDNet [29] 71.4 66.7 87.7 60.6 77.8 50.7 28.9 87.1 74.3 56.8 76.3 66.9
+SAFDNet [28] 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
+LION [14] 72.1 68.0 87.9 64.9 77.6 44.4 28.5 89.6 75.6 59.4 80.8 71.6
+FSHNet[13] 71.7 68.1 88.7 61.4 79.3 47.8 26.3 89.3 76.7 60.5 78.6 72.3
+UniMamba [11] 72.6 68.5 88.7 64.7 79.7 47.9 28.7 89.7 74.6 59.1 79.5 72.3
 VDM-mamba(Ours) 72.5 68.1 88.5 59.5 78.8 43.2 30.6 90.2 75.2 61.1 81.4 72.5
-VDM-OD-mamba(Ours) 72.9 68.5 88.7 65.7 78.9 43.8 29.7 90.2 75.7 60.9 81.1 70.3
+VDM-OD (Ours) 72.9 68.5 88.7 65.7 78.9 43.8 29.7 90.2 75.7 60.9 81.1 70.3
 Performances on the test set
-TransFusion-LiDAR 70.2 65.5 86.2 56.7 66.3 58.8 28.2 86.1 68.3 44.2 82.0 78.2
-DSVT 72.7 68.4 86.8 58.4 67.3 63.1 37.1 88.0 73.0 47.2 84.9 78.4
-HEDNet 72.0 67.7 87.1 56.5 70.4 63.5 33.6 87.9 70.4 44.8 85.1 78.1
-VPF 72.7 67.0 85.8 55.1 63.5 62.1 33.3 87.6 72.5 48.6 82.9 78.2
-SAFDNet 72.3 68.3 87.3 57.3 68.0 63.7 37.3 89.0 71.1 44.8 84.9 79.5
-LION 73.9 69.8 87.2 61.1 68.9 65.0 36.3 90.0 74.0 49.2 87.3 79.5
+TransFusion-LiDAR [1] 70.2 65.5 86.2 56.7 66.3 58.8 28.2 86.1 68.3 44.2 82.0 78.2
+DSVT [24] 72.7 68.4 86.8 58.4 67.3 63.1 37.1 88.0 73.0 47.2 84.9 78.4
+HEDNet [29] 72.0 67.7 87.1 56.5 70.4 63.5 33.6 87.9 70.4 44.8 85.1 78.1
+VPF [10] 72.7 67.0 85.8 55.1 63.5 62.1 33.3 87.6 72.5 48.6 82.9 78.2
+SAFDNet [28] 72.3 68.3 87.3 57.3 68.0 63.7 37.3 89.0 71.1 44.8 84.9 79.5
+LION [14] 73.9 69.8 87.2 61.1 68.9 65.0 36.3 90.0 74.0 49.2 87.3 79.5
+UniMamba[11] 74.0 70.2 87.9 60.4 70.9 65.9 36.7 90.5 73.5 49.5 86.9 79.4
 VDM-mamba(Ours) 73.7 70.0 88.0 60.1 67.6 66.0 35.1 90.9 72.2 50.9 89.2 80.0
-Table 3: Performances on the nuScenes validation and test set.
+VDM-OD (Ours) 73.9 70.5 88.3 62.2 70.7 65.3 35.0 90.6 74.4 51.1 88.3 78.8
+TABLE III: Performances on the nuScenes validation and test set.
 ‘T.L.’, ‘C.V.’, ‘Ped.’, ‘M.T.’, ‘T.C.’, and ’B.R.’ are short for trailer, construction vehicle, pedestrian, motor, traffic cone, and barrier, respectively.
-VDM-OD-Mamba denotes a variant where the Only-Diffusion structure is used within the VDM framework.
+VDM-OD denotes a variant where the Only-Densification structure is used within the VDM framework.
 Method
@@
 A-Bus
-Shool Bus
+School Bus
 Truck Cab
@@
 W-Rider
-CenterPoint 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1 22.1 0.0 3.9 0.5 10.9 4.2
-HEDNet 37.1 78.2 47.7 67.6 46.4 45.9 56.9 67.0 48.7 46.5 58.2 47.5 23.3 40.9 27.5 46.8 27.9 20.6 6.9 27.2 38.7 21.6 0.0 30.7 9.5 28.5 8.7
-VoxelNeXt 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
-FSDV2 37.6 77.0 47.6 70.5 43.6 41.5 53.9 58.5 56.8 39.0 60.7 49.4 28.4 41.9 30.2 44.9 33.4 16.6 7.3 32.5 45.9 24.0 1.0 12.6 17.1 26.3 17.2
-SAFDNet 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
-LION 41.5 75.1 43.6 73.9 53.9 45.1 66.4 74.7 61.3 48.7 65.1 56.2 21.7 42.7 25.3 58.4 28.9 23.6 8.3 49.5 47.3 19.0 0.0 31.4 8.7 37.6 11.8
-UniMamba 42.0 78.9 47.9 74.3 51.8 46.8 67.8 76.9 55.8 51.7 62.8 52.2 30.2 44.6 24.6 28.1 59.4 32.2 23.2 6.7 41.5 48.5 0.0 26.4 8.1 36.4 13.7
+CenterPoint [27] 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1 22.1 0.0 3.9 0.5 10.9 4.2
+HEDNet [29] 37.1 78.2 47.7 67.6 46.4 45.9 56.9 67.0 48.7 46.5 58.2 47.5 23.3 40.9 27.5 46.8 27.9 20.6 6.9 27.2 38.7 21.6 0.0 30.7 9.5 28.5 8.7
+VoxelNeXt [5] 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
+FSDV2 [7] 37.6 77.0 47.6 70.5 43.6 41.5 53.9 58.5 56.8 39.0 60.7 49.4 28.4 41.9 30.2 44.9 33.4 16.6 7.3 32.5 45.9 24.0 1.0 12.6 17.1 26.3 17.2
+SAFDNet [28] 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
+LION [14] 41.5 75.1 43.6 73.9 53.9 45.1 66.4 74.7 61.3 48.7 65.1 56.2 21.7 42.7 25.3 58.4 28.9 23.6 8.3 49.5 47.3 19.0 0.0 31.4 8.7 37.6 11.8
+FSHNet[13] 40.2 77.4 48.3 72.9 50.0 47.2 63.4 69.9 56.1 43.8 62.5 53.7 31.8 44.8 28.9 56.6 32.0 23.0 8.7 34.5 41.0
+UniMamba [11] 42.0 78.9 47.9 74.3 51.8 46.8 67.8 76.9 55.8 51.7 62.8 52.2 30.2 44.6 24.6 28.1 59.4 32.2 23.2 6.7 41.5 48.5 0.0 26.4 8.1 36.4 13.7
 VDM (Ours) 42.3 78.2 48.2 75.2 53.2 45.7 68.2 77.4 60.3 47.2 67.0 58.0 24.9 43.0 24.9 60.1 31.4 25.1 7.8 42.5 46.3 20.9 0.0 29.7 10.3 39.8 13.6
 VDM-OD (Ours) 42.6 79.0 49.5 75.9 54.5 45.5 69.9 77.9 56.9 47.6 66.6 59.2 24.0 44.2 24.7 60.9 33.2 24.9 7.5 39.5 51.3 20.1 0.0 33.7 7.8 40.1 12.5
-Table 4: Comparison with prior methods on the Argoverse2 validation set.
+TABLE IV: Comparison with prior methods on the Argoverse2 validation set.
 ‘Vehicle’, ‘C-Barrel’, ‘MPC-Sign’, ‘A-Bus’, ‘C-Cone’, ‘V-Trailer’, ‘MBT’, ‘W-Device’, and ‘W-Rider’ are abbreviations for regular vehicle, construction barrel, mobile pedestrian crossing sign, articulated bus, construction cone, vehicular trailer, message board trailer, wheeled device, and wheeled rider, respectively.
-VDM and VDM-OD refer to the proposed VDM-Mamba and VDM-OD-Mamba variants.
-Implementation Details
+VDM and VDM-OD refer to the proposed VDM-Mamba and VDM-OD variants.
+IV Experiments
+IV-A Datasets and Evaluation Metrics
+Waymo Open Dataset.
+The Waymo Open Dataset (WOD) [21] includes 1,150 scenes, divided into 798 for training, 202 for validation, and 150 for testing.
+Each scene covers a perception range of $150\,\text{m}\times 150\,\text{m}$.
+For evaluation, WOD adopts 3D mean Average Precision (mAP) and mAP weighted by heading accuracy (mAPH).
+Each metric includes two difficulty levels: L1 for objects with more than five points, and L2 for bounding boxes containing one to five points.
+nuScenes Dataset.
+The nuScenes dataset [3] contains 1,000 scenes, with 750 for training, 150 for validation, and 150 for testing.
+It uses mean Average Precision (mAP) and nuScenes Detection Score (NDS) as evaluation metrics.
+The dataset covers 10 annotated object classes, representing most common categories in traffic scenarios.
+Argoverse 2 Dataset.
+The Argoverse 2 dataset [4] comprises 1,000 sequences, with 700 for training, 150 for validation, and 150 for testing.
+The perception range extends up to 200 meters.
+It uses mean Average Precision (mAP) as the evaluation metric.
+ONCE Dataset.
+The ONCE dataset [16] consists of 5,000, 3,000, and 8,000 frames for training, validation, and testing respectively.
+It defines three detection classes: Vehicle, Pedestrian, and Cyclist.
+The final performance is evaluated using mean Average Precision (mAP) across all three classes.
+IV-B Implementation Details
 Network Architecture.
-We evaluate the proposed VDM module by integrating it with both the SSMs-based LION (Liu et al.
-2024) and the Transformer-based DSVT (Wang et al.
-2023) models.
+We evaluate the proposed VDM module by integrating it with both the SSMs-based LION [14] and the Transformer-based DSVT [24] models.
 We use voxel sizes of $(0.075\,\text{m},0.075\,\text{m},0.25\,\text{m})$, $(0.08\,\text{m},0.08\,\text{m},0.1875\,\text{m})$, $(0.1\,\text{m},0.1\,\text{m},0.25\,\text{m})$, and $(0.1\,\text{m},0.1\,\text{m},0.25\,\text{m})$ for the nuScenes, Waymo, Argoverse 2, and ONCE datasets respectively.
 The number of voxel feature channels is set to 64.
+Specifically for the nuScenes dataset, we incorporate a Sparse Residual Module to enhance spatial feature interactions among voxels.
 In the VDM module, the SubM3D (including SRB) layers use a kernel size of 3.
@@
 For LION and DSVT, we follow the default parameter settings from their original papers without modification.
-Additionally, we conduct an only-diffusion experiment, in which the voxel size along the $x$ and $y$ axes is scaled by a factor of 4, while the $z$ axis remains unchanged.
+Additionally, we conduct an only-densification experiment, in which the voxel size along the $x$ and $y$ axes is scaled by a factor of 4, while the $z$ axis remains unchanged.
 Training Process.
@@
 On the Waymo Open Dataset, the VDM-SSMs configuration is trained for 24 epochs with a batch size of 16, while the VDM-DSVT configuration uses 24 epochs with a batch size of 8.
-For the remaining three datasets, we adopt the same training strategy as that used in the LION (Liu et al.
-2024) baseline.
-Main Results
+For the remaining three datasets, we adopt the same training strategy as that used in the LION [14] baseline.
+IV-C Main Results
 To verify the effectiveness of the proposed module, we design two models: VDM-SSMs based on LION and VDM-Transformer based on DSVT.
@@
 Results on WOD.
-As shown in Table 2, we incorporate the VDM module into both Mamba-based and Transformer-based models.
-Among them, VDM-Mamba achieves a state-of-the-art (SOTA) result of 74.8 mAPH (L2), outperforming the corresponding baseline LION model (without VDM) by +0.7 mAPH, and also surpassing the recent UniMamba (Jin et al.
-2025).
+As shown in Table II, we incorporate the VDM module into both Mamba-based and Transformer-based models.
+Among them, VDM-Mamba achieves a state-of-the-art (SOTA) result of 74.8 mAPH (L2), outperforming the corresponding baseline LION model (without VDM) by +0.8 mAPH, and also surpassing the recent UniMamba [11].
 In addition, VDM-DSVT also demonstrates performance gains, achieving a +1.2 mAPH (L2) improvement over the original DSVT, which validates the general applicability of the VDM module to both SSM-based and Transformer-based detectors.
-When only the diffusion mechanism is retained in VDM (i.e., VDM-OD-Mamba), the model still achieves the best results on the Vehicle and Cyclist categories, highlighting the importance of the diffusion component in the overall design.
+When only the densification mechanism is retained in VDM (i.e., VDM-OD), the model still achieves the best results on the Vehicle and Cyclist categories, highlighting the importance of the densification component in the overall design.
 Moreover, VDM-Mamba achieves the best performance in the Pedestrian category, suggesting that the fine-grained aggregation ability introduced by VDM is particularly beneficial for small object detection.
@@
 overall 0-30m 30-50m 50m+ overall 0-30m 30-50m 50m+ overall 0-30m 30-50m 50m+
-PointRCNN 52.1 74.5 40.9 16.8 4.3 6.2 2.4 0.9 29.8 46.0 20.9 5.5 28.7
-PointPillars 68.6 80.9 62.1 47.0 17.6 19.7 15.2 10.2 46.8 58.3 40.3 25.9 44.3
-SECOND 71.2 84.0 63.0 47.3 26.4 29.3 24.1 18.1 58.0 70.0 52.4 34.6 51.9
-PV-RCNN 77.8 89.4 72.6 58.6 23.5 25.6 22.8 17.3 59.4 71.7 52.6 36.2 53.6
-CenterPoint 66.8 80.1 59.6 43.4 49.9 56.2 42.6 26.3 63.5 74.3 57.9 41.5 60.1
-PointPainting 66.2 80.3 59.8 42.3 44.8 52.6 36.6 22.5 62.3 73.6 57.2 40.4 57.8
-LION 78.2 89.1 72.6 57.5 53.2 62.4 44.0 24.5 68.5 79.2 63.2 47.1 66.6
-VDM-mamba 78.6 89.1 74.3 59.3 54.2 63.3 43.2 26.1 69.9 79.4 64.9 49.8 67.6
-Table 5: Comparison with previous methods on ONCE validation set.
+PointRCNN [20] 52.1 74.5 40.9 16.8 4.3 6.2 2.4 0.9 29.8 46.0 20.9 5.5 28.7
+PointPillars [12] 68.6 80.9 62.1 47.0 17.6 19.7 15.2 10.2 46.8 58.3 40.3 25.9 44.3
+SECOND [25] 71.2 84.0 63.0 47.3 26.4 29.3 24.1 18.1 58.0 70.0 52.4 34.6 51.9
+PV-RCNN [18] 77.8 89.4 72.6 58.6 23.5 25.6 22.8 17.3 59.4 71.7 52.6 36.2 53.6
+CenterPoint [27] 66.8 80.1 59.6 43.4 49.9 56.2 42.6 26.3 63.5 74.3 57.9 41.5 60.1
+PointPainting [23] 66.2 80.3 59.8 42.3 44.8 52.6 36.6 22.5 62.3 73.6 57.2 40.4 57.8
+LION [14] 78.2 89.1 72.6 57.5 53.2 62.4 44.0 24.5 68.5 79.2 63.2 47.1 66.6
+VDM-mamba (Ours) 78.6 89.1 74.3 59.3 54.2 63.3 43.2 26.1 69.9 79.4 64.9 49.8 67.6
+VDM-OD (Ours) 78.5 88.8 73.1 59.7 51.0 59.2 42.3 24.5 68.7 79.1 64.1 48.9 66.1
+TABLE V: Comparison with previous methods on ONCE validation set.
 Results on nuScenes.
-To further verify the effectiveness of the proposed VDM module, we conduct experiments on both the nuScenes validation and test sets.
-As shown in Table 3, VDM-mamba achieves 68.1 mAP and 72.5 NDS on the validation set, and 70.0 mAP and 73.7 NDS on the test set.
-Although VDM-mamba slightly underperforms UniMamba in terms of overall detection accuracy, it achieves the best results in the pedestrian and bicycle categories on the nuScenes validation set.
-These categories are particularly important for autonomous driving applications, demonstrating that our proposed method offers clear advantages in detecting small and vulnerable objects.
-Furthermore, VDM-OD-Mamba achieves a state-of-the-art 72.9 NDS, suggesting that the diffusion-only structure plays a critical role in improving detection performance.
+To further verify the effectiveness of the proposed VDM, we conduct experiments on both the nuScenes validation and test sets.
+As shown in Table III, on the validation set, VDM-Mamba achieves 68.1 mAP and 72.5 NDS, while the VDM-OD variant achieves even higher performance with 68.5 mAP and 72.9 NDS.
+Most notably, on the test set, VDM-OD achieves 70.5 mAP, surpassing the previous state-of-the-art UniMamba (70.2 mAP) and LION (69.8 mAP), while maintaining a competitive 73.9 NDS.
+Detailed analysis reveals that VDM-Mamba excels in detecting vulnerable road users, achieving the best results in the Pedestrian and Bicycle categories on the validation set.
+This demonstrates that the fine-grained aggregation component in the full VDM offers clear advantages for small object localization.
+However, the superior overall performance of VDM-OD (70.5 mAP on Test) suggests that the densification-only structure plays the most critical role in mitigating feature sparsity and boosting general detection accuracy.
 Results on Argoverse V2.
-As shown in Table 4, our proposed models, VDM-Mamba and VDM-OD-Mamba, achieve state-of-the-art detection performance on the Argoverse 2 validation set, reaching 42.3 mAP and 42.6 mAP, respectively.
-Compared with the strong baseline LION (Liu et al.
-2024), VDM-Mamba yields a notable improvement of 0.8 mAP.
-Notably, VDM-OD-Mamba, which retains only the diffusion structure within the VDM module, still surpasses all previous approaches, highlighting the effectiveness of the diffusion mechanism.
+As shown in Table IV, our proposed models, VDM-Mamba and VDM-OD, achieve state-of-the-art detection performance on the Argoverse 2 validation set, reaching 42.3 mAP and 42.6 mAP, respectively.
+Compared with the strong baseline LION [14], VDM-Mamba yields a notable improvement of 0.8 mAP.
+Notably, VDM-OD, which retains only the densification structure within the VDM module, still surpasses all previous approaches, highlighting the effectiveness of the densification mechanism.
 These results further demonstrate the strong representational capacity and scalability of the VDM module in large-scale point cloud scenarios.
-Results on ONCE.
-VDM-Mamba also demonstrates superior performance on the ONCE dataset.
-It achieves detection accuracies of 78.6 mAP, 54.2 mAP, and 69.9 mAP on the Vehicle, Pedestrian, and Cyclist categories, respectively.
-The overall detection accuracy across all categories reaches 67.6 mAP, representing a +1.0 mAP improvement over the Mamba-based LION detector.
-These results establish VDM-Mamba as the new state-of-the-art (SOTA) on the ONCE benchmark.
-Ablation Study
-Diffusion Waymo(3D AP/APH)L2
-Vehicle Pedestrian Cyclist
-No(LION) 68.8/68.3 77.3/72.6 74.7/73.8
-Yes(VDM-OD) 70.3/69.9 78.2/73.8 76.0/75.0
-Table 6: Detection results of the voxel diffusion module on the Waymo validation set (20% data).
-”No” indicates that the original voxels are not diffused, and the detection results are entirely determined by LION(Liu et al.
-2024) ”Yes” indicates that we only adopt the diffusion module from VDM, i.e., using sparse 3D spconv with a kernel size of 3 and a stride of 1.
-Ablation Study on Voxel Diffusion.
-To assess the contribution of voxel diffusion to detection performance, we conduct ablation experiments on the Waymo validation set using a fixed voxel size of [0.32, 0.32, 0.1875].
-We evaluate the model both with and without voxel diffusion.
-As shown in Table 6, introducing voxel diffusion improves the VDM-OD model by 1.3 mAPH (L2), underscoring its critical role in sequential detection frameworks.
-Moreover, on the nuScenes (Table 3) and Argoverse 2 (Table 4) datasets, we observe that under the Only-diffusion setting, the model still achieves higher detection accuracy, further validating the significant impact of voxel diffusion.
 Fine-grained ONCE (Overall) Waymo (Cyclist)
@@
 Fine (VDM) 78.6 54.2 69.9 67.6 79.1 78.1 76.2 75.2
-Table 7: Ablation study on the ONCE and Waymo (20% data) dataset for fine-grained feature aggregation.
+TABLE VI: Ablation study on the ONCE and Waymo (20% data) dataset for fine-grained feature aggregation.
 Base denotes the detection results obtained by combining the VDM and Mamba network without fine voxel partitioning.
 Fine represents the results after applying fine-grained voxel partitioning.
+Method Speed (s)
+Waymo nuScenes Argoverse2 ONCE
+LION 0.206 0.213 0.185 0.149
+VDM 0.357 0.291 0.320 0.228
+TABLE VII: Runtime comparison on different datasets. The inference speed is measured in seconds (s) on a single NVIDIA 3090 GPU with a batch size of 1.
+Densification Waymo(3D AP/APH)L2
+Vehicle Pedestrian Cyclist
+No(LION) 68.8/68.3 77.3/72.6 74.7/73.8
+Yes(VDM-OD) 70.3/69.9 78.2/73.8 76.0/75.0
+TABLE VIII: Detection results of the voxel densification module on the Waymo validation set (20% data).
+“No” indicates that the original voxels are not diffused, and the detection results are entirely determined by LION [14].
+“Yes” indicates that we only adopt the densification module from VDM, i.e., using sparse 3D spconv with a kernel size of 3 and a stride of 1.
+Results on ONCE.
+VDM-Mamba demonstrates superior performance on the ONCE dataset (Table V).
+It achieves detection accuracies of 78.6 mAP, 54.2 mAP, and 69.9 mAP on the Vehicle, Pedestrian, and Cyclist categories, respectively.
+The overall detection accuracy reaches 67.6 mAP, representing a +1.0 mAP improvement over the baseline LION detector, establishing VDM-Mamba as the new state-of-the-art (SOTA).
+In contrast to the nuScenes results, the VDM-OD variant (utilizing only densification) achieves 66.1 mAP, which is lower than the full VDM framework.
+This performance gap highlights the critical contribution of the fine-grained feature aggregation component on this dataset.
+While voxel expansion addresses sparsity, the aggregation module captures detailed local geometric cues that prove essential for the ONCE benchmarks, validating the necessity of our dual-path design (densification + aggregation) for robust detection across diverse scenarios.
+Qualitative Visualization.
+To visually illustrate the advanced capabilities of our proposed model (VDM), we present a qualitative comparison of its detection performance against LION [14] on the Waymo validation set.
+As depicted in Figure 5, the annotations marked by yellow arrows highlight several advantages of our method.
+Firstly, the bounding boxes predicted by VDM exhibit a higher degree of alignment with the ground truth annotations.
+Secondly, VDM is capable of correctly detecting objects (true positives) that are entirely missed by LION.
+Finally, our model demonstrates a lower false positive rate, indicating superior precision.
+Runtime Comparison.
+Table VII compares the inference time of our VDM with LION [14].
+While VDM’s voxel densification approach increases latency by processing more voxels, it offers a favorable accuracy-efficiency trade-off.
+On the nuScenes and ONCE datasets, latency increases by only 0.08s for improvements of 0.4 in NDS and 1.0 in mAP.
+For long-range detection on datasets like Waymo and Argoverse 2, the more substantial latency is justified by significant performance gains of 0.8 mAPH (L2) and 0.8 mAP, respectively.
+IV-D Ablation Study
+Figure 5: A comparison of point cloud detection results between VDM and LION on the Waymo validation set.
+The blue, green, and red boxes represent human annotations, VDM predictions, and LION predictions, respectively.
+Yellow arrows on the figures highlight the differences in detection results.
+In the point cloud visualization process, the score threshold is set to 0.3, and the non-maximum suppression (NMS) threshold is set to 0.1.
+Ablation Study on Voxel Densification.
+To assess the contribution of voxel densification to detection performance, we conduct ablation experiments on the Waymo validation set using a fixed voxel size of [0.32, 0.32, 0.1875].
+We evaluate the model both with and without the densification mechanism.
+As shown in Table VIII, introducing voxel expansion improves the VDM-OD model by approximately 1.3 mAPH (L2), underscoring its critical role in sequential detection frameworks.
+Moreover, on the nuScenes (Table III) and Argoverse 2 (Table IV) datasets, we observe that under the Only-Densification setting, the model achieves impressive detection accuracy, further validating the significant impact of pre-serialization voxel expansion.
 Ablation on Fine-grained Feature Aggregation.
-To further evaluate the fine-grained feature aggregation capability of VDM under more detailed voxel partitioning, we conducted experiments by refining the voxel resolution.
-Specifically, we reduced the voxel size along the x and y axes to one-fourth of the original size, while keeping the resolution along the z-axis unchanged, with voxel diffusion still enabled.
-As shown in Table 2, under this fine-grained aggregation setting, the model achieves a 0.5 APH (L2) improvement in detecting the Pedestrian category, indicating enhanced localization and recognition of small-scale objects.
-Furthermore, as reported in Table 7, on the ONCE dataset, the VDM-OD model exhibits inferior performance across all three categories when only voxel diffusion is applied.
-With the integration of fine-grained feature aggregation, the model achieves a 1.5 mAP gain, demonstrating the effectiveness of finer voxel partitioning in enhancing detection performance.
-When trained with 20% of the Waymo validation set, the VDM model also achieves higher detection accuracy for the Cyclist category.
-Conclusion
-We propose the Voxel Diffusion Module (VDM), which consists of a sparse 3D convolutional network, sparse residual blocks, and submanifold sparse convolution (SubM3D).
-This module diffuses voxelized point clouds via sparse 3D convolutions to increase the density of foreground voxels for more accurate object localization, while simultaneously performing spatial pre-aggregation of voxel features to provide spatial context for the subsequent serialized voxel sequences.
-To validate the effectiveness of our approach, we evaluate two model variants: VDM-LION (based on Mamba) and VDM-DSVT (based on Transformer).
-Experimental results demonstrate that VDM consistently improves performance in both SSMs-based and Transformer-based architectures.
-Notably, VDM-LION (SSMs) achieves state-of-the-art results on the Waymo, Argoverse 2, and ONCE datasets.
+To further evaluate the fine-grained feature aggregation capability of VDM, we conducted experiments by refining the voxel resolution.
+Specifically, we reduced the voxel size along the x and y axes to one-fourth of the original size, while keeping the resolution along the z-axis unchanged, with voxel expansion still enabled.
+As reported in Table VI, on the ONCE dataset, the Base model (VDM-OD) exhibits inferior performance across all three categories when only voxel expansion is applied.
+With the integration of fine-grained feature aggregation (Fine), the model achieves a +1.5 mAP gain, demonstrating the overall effectiveness of finer voxel partitioning.
+To delve deeper into these improvements, we analyze the performance per category.
+It is worth noting that while the simplified VDM-OD (densification-only) variant achieves impressive performance on rigid, large-scale objects such as Vehicles (e.g., 70.5 mAP on nuScenes test set), the full VDM framework demonstrates superior capability in detecting smaller, non-rigid objects.
+For instance, on the ONCE validation set, the full VDM outperforms VDM-OD by a significant margin on the Pedestrian category (+3.2 mAP in Table V).
+We attribute this phenomenon to the inherent geometric differences between categories.
+Large rigid objects (Vehicles) benefit primarily from increased voxel density to recover structural completeness, which VDM-OD provides efficiently.
+In contrast, small and sparse objects (Pedestrians) possess fewer points and more complex local geometries.
+The fine-grained aggregation mechanism in the full VDM is essential for capturing these subtle local contexts, which might be overlooked by simple densification.
+Therefore, although VDM-OD offers a strong baseline for vehicles, the full VDM provides a more balanced and robust solution, particularly for vulnerable road users (VRUs) critical to autonomous driving safety.
+V Conclusion
+In this paper, we propose the Voxel Densification Module (VDM), a novel pre-serialization stem designed to bridge the spatial-serial gap in existing 3D detectors.
+By leveraging sparse 3D convolutions, VDM explicitly expands the foreground voxel set to address the sparsity limitation inherent in serialized models, while simultaneously performing fine-grained spatial aggregation to enrich local context.
+We validate the universality of our approach by integrating VDM into both SSM-based (LION) and Transformer-based (DSVT) architectures.
+Extensive experiments across four large-scale benchmarks demonstrate that VDM consistently boosts detection performance.
+Our analysis reveals that while pre-serialization densification serves as the primary driver for mitigating feature sparsity, the fine-grained aggregation mechanism provides specific gains in scenarios requiring richer local geometric details.
+These results confirm that explicit voxel densification is a generic and effective strategy for advancing point cloud object detection.